### v2

# FarsInstruct: Empowering Large Language Models for Persian Instruction Understanding

Abstract
Instruction-tuned large language models, such as T0, have demonstrated remarkable capabilities in following instructions across various domains. However, their proficiency remains notably deficient in many low-resource languages. To address this challenge, we introduce FarsInstruct: a comprehensive instruction dataset designed to enhance the instruction-following ability of large language models specifically for the Persian language—a significant yet underrepresented language globally. FarsInstruct encompasses a wide range of task types and datasets, each containing a mix of straightforward to complex manual written instructions, as well as translations from Public Pool of Prompts, ensuring a rich linguistic and cultural representation. Furthermore, we introduce Co-CoLA, a framework designed to enhance the multi-task adaptability of LoRA-tuned models. Through extensive experimental analyses, our study showcases the effectiveness of FarsInstruct dataset coupled with training by Co-CoLA framework [^1] https://github.com/Hojjat-Mokhtarabadi/FarsInstruct, in improving the performance of large language models within the Persian context. As of the current writing, FarsInstruct comprises more than 200 templates across 21 distinct datasets, and we intend to update it consistently, thus augmenting its applicability.
Keywords: Instruction-tuned LLMs, Low-resource languages, Parameter efficient fine-tuning
FarsInstruct: Empowering Large Language Models
for Persian Instruction Understanding
Hojjat Mokhtarabadithanks: All the authors are members of the Department of Computer Engineering. h.mokhtarabadi@ec.iut.ac.ir Isfahan University of Technology                        Ziba Zamanithanks: Zibazamani@eng.uk.ac.ir Shahid Bahonar University of Kerman
Abbas Maazallahithanks: a.maazallahi@ut.ac.ir University of Tehran                        Hossein Manshaeithanks: manshaei@iut.ac.ir Isfahan University of Technology
1 Introduction
The modern era of artificial intelligence is marked by numerous breakthroughs, among which is the rise of large language models (LLMs). These models, such as PaLM (Chowdhery et al., 2022), GPT4 (OpenAI et al., 2024), and Llama2 (Touvron et al., 2023) with continuous scaling of their parameters and training data, are known to exhibit emergent properties. Wei et al. (2022a) considers an ability to be emergent if it is not present in smaller models but is present in larger models. This is an unpredictable phenomenon that can not be predicted simply by extrapolating the performance of smaller models. One such ability is instruction-following, which enables models to execute unseen natural language processing (NLP) tasks from reading instructions provided within the input text. Previously, the capability for instruction-following was primarily attributed to the scale of these models. However, recent studies have demonstrated that instruction-following does not exclusively rely on the large size of language models (Sanh et al., 2022). By instruction-tuning on a collection of instructional NLP tasks, smaller language models can learn to follow prompts. This approach has proven to be remarkably efficient, allowing these smaller models to perform competitively and, in some specific tasks, even outperform their larger counterparts (Sanh et al., 2022; Wei et al., 2021; Wang et al., 2022). Instruction-tuning emerges as a vital technique in the evolution of language models, involving training a model on a wide range of tasks described through natural language instructions. This method diverges from traditional task-specific fine-tuning, offering a more generalized and versatile approach to model training, thus contributing significantly to the advancement of LLMs.
Despite the steady progress of instruction-tuned language models, they still struggle to accurately grasp the subtleties of low-resourced languages due to the scarcity of prompted data and challenges inherent in translating English datasets into other languages (Naous et al., 2024; Ramesh et al., 2023; Vanmassenhove et al., 2021). While efforts have been made to compile extensive multilingual instruction-following datasets, gaps remain in creating diverse and complex prompts for languages like Persian. For example, the SuperNaturalInstructions benchmark (Wang et al., 2022), encompassing various task types across 55 languages, contains merely 2.1% of Persian content. Similarly, the Aya Dataset (Singh et al., 2024), a human-curated effort to enhance AI’s instruction-following abilities across 65 languages, includes 1% of Persian content. This underscores the disparity in the diversity and quantity of the Persian Language tasks compared to other languages.
In this study, we propose FarsInstruct, a comprehensive prompted dataset tailored to the Persian language. It comprises a mixture of manually written instructions ranging from basic to proficient language levels, as well as translations from Public Pool of Prompts (P3) (Sanh et al., 2022) which is a collection of prompted English datasets. In particular, we created more than 200 prompt templates (roughly 10 templates for each of the 21 unique public datasets) that we selected from a variety of sources. These datasets collectively cover ten different task categories: Text Summarization, Textual Entailment, Text Classification, Sentiment Analysis, Word Sense Disambiguation, Query Paraphrasing, Question Answering, Reading Comprehension, Named Entity Recognition (NER), and Translation. Figure 1 depicts an instance of a prompt within our dataset, and a detailed overview of FarsInsturct dataset is provided in Section 3.
Figure 1: An example of the prompts utilized in the training process. The Persian version of the prompt is employed for training purposes, while the translated English version is provided to enhance comprehension. The instruction component is highlighted in blue, the data field is marked in orange, and the target answer is indicated in gray. In Appendix C, this example is shown in PromptSource environment.
Additionally, in order to facilitate the multi-task adaptation of our model and mitigate the problem of catastrophic forgetting, we introduce Co-CoLA, an integration of CoLA (Xia et al., 2024) with rehearsal training (Kirkpatrick et al., 2017). More specifically, we adopt an iterative optimization framework that merges learned low-rank matrices into the model parameters and reinitializes optimization for new LoRA modules. At each iteration, we involve retraining a subset of data from previously learned tasks and mixing it with the current task’s data during training. With this periodic revisiting of earlier tasks, the model maintains performance on both old and new tasks while preserving computational efficiency. Section 4 presents an in-depth explanation of Co-CoLA method.
FarsInstruct is publicly available and open-source and we are committed to enhancing it by continually expanding our dataset with a broader range of tasks, instruction entries, and modalities. We hope this dataset fills the critical gap and serves as a valuable resource to the NLP community.
2 Related work
Instruction-tuning: In the landscape of AI, the capabilities of LLMs have expanded far beyond mere text processing. These sophisticated models are now being fine-tuned in a practice known as instruction-tuning, where models are trained with specific input-output pairs drawn from a wide array of data sources. This technique enables a pre-trained LLM to produce tailored outputs based on given inputs, enhancing its versatility and effectiveness. FLAN (Wei et al., 2021) and T0 (Sanh et al., 2022) pioneered the exploration of instruction-tuned language models, each contributing significantly to the field. FLAN (Wei et al., 2021) adapted a 137-billion parameter pre-trained model, refining it with over 60 NLP datasets using natural language instructions. On the other hand, T0 (Sanh et al., 2022) applied instruction tuning to various T5 models across 2073 prompts from 177 datasets. SuperNaturalInstruction (Wang et al., 2022) further advanced the field by assembling a comprehensive benchmark featuring 1,616 expert-written NLP tasks, covering 76 unique task types, and extending support to multiple languages. xP3 (Muennighoff et al., 2022) expanded on P3’s groundwork (Sanh et al., 2022) by including content from 46 languages, adding new tasks like Translation and Program Synthesis that P3 had not tackled. In a similar expansive effort, Aya (Singh et al., 2024) emerged as a significant multilingual project, featuring an impressive collection of 513 million instances across 114 languages, achieved through a collaborative research effort that involved fluent speakers from around the world to compile and complete instructional content. Our dataset distinguishes itself from these collections in its depth and adaptability, especially with the inclusion of more challenging Persian tasks, offering a high level of detail not found in many multilingual efforts. While most such projects primarily use machine translations and cover a narrow range of tasks, our dataset presents a wide array of culturally and linguistically rich tasks.
Parameter effecient fine-tuning: Conventional full-parameter fine-tuning becomes computationally impractical as the model size and the number of downstream tasks increase. To address this challenge, recent advancements in parameter-efficient fine-tuning methods suggest training only a small portion of parameters while keeping the majority of pre-trained model parameters unchanged. One of the most widely used paradigms in parameter-efficient fine turning is Low-Rank Adaptation (LoRA) (Hu et al., 2021). LoRA only modifies a small, low-rank portion of the model’s weights. This is achieved by adding low-rank matrices to the model’s weights during training. Despite the significant computational advantage of LoRA, it falls short in multi-task adaptation, and also Kalajdzievski (2024) showed that PEFT strategies, such as LoRA, are still susceptible to catastrophic forgetting. MultiLoRA (Wang et al., 2023) addresses the limitations of LoRA by reducing the dominance of top singular vectors, horizontally scaling LoRA modules, and altering the initialization of adaptation matrices, which leads to improved performance across multiple tasks with minimal additional parameters. MixLoRA (Li et al., 2024) introduces multiple LoRA-based experts within a frozen pre-trained model using a top-k routing strategy to efficiently distribute tasks, independently configure attention layer adapters, and apply auxiliary load balance loss, significantly enhancing performance while reducing GPU memory consumption and training latency. Additionally, CoLA(Xia et al., 2024) introduces an iterative optimization framework designed to improve the fine-tuning of LLMs by employing multiple iterations of LoRA. In this paper, we design Co-CoLA to address the issue of catastrophic forgetting, while ensuring an effective multi-task adaption.
3 FarsInstruct Dataset
With about 130 million [^2] https://en.wikipedia.org/wiki/Persian_language speakers, Persian — also referred to as Farsi in Iran — is an important language in the Middle East and Central Asia. FarsInstruct represents a project to provide a comprehensive public prompted dataset for the Persian community. As of this writing, FarsInstruct has more than 200 carefully designed and created prompt templates for 21 already-published public datasets and some translations from existing prompted datasets. Unlike multilingual collections focusing on common tasks such as Text Summarization and Question Answering, FarsInstruct introduces more innovative and challenging tasks, including Named Entity Recognition and Word Sense Disambiguation. The creation procedure, statistics, task augmentation, and quality of the dataset are covered in detail in the following subsections. Additional illustrations and tables are provided in the Appendix BC.
3.1 Dataset Construction
The development of FarsInstruct entailed transforming Persian NLP datasets into their prompted format, described in plain language. This process involved a combination of manual ideation, during which our team meticulously brainstormed and refined prompt templates, along with invaluable insights from Persian language instructors. For datasets with multiple data fields, prompts were crafted to interrelate these fields, as elaborated in Section 3.2. Additionally, synonyms were employed to diversify the instructions within the prompts and reduce repetition. Each prompt template falls into one of two classes: categorization or generation. Categorization prompts guide the model in classifying text into predefined categories from dataset labels or identified through dataset analysis. In contrast, generation prompts require the model to produce full-length text, such as summarizing longer texts or answering questions based on the provided information. These instructions also include scenarios where the model needs to generate missing content from partial text inputs.
Figure 2: Detailed depiction of 11 task types utilized in our dataset. Each box within the figure lists the specific datasets associated with the respective task type. Datasets designated for training are highlighted in blue, and those reserved for testing are marked in orange. Additionally, manual datasets, which have been specifically curated and prompted by our team, are enclosed with solid borders. In contrast, datasets that have been translated from English to Persian are enclosed with dashed borders.
To efficiently create a large collection of prompts, we primarily utilized PromptSource (Bach et al., 2022), an open-source tool designed for creating, sharing, and managing prompts for NLP tasks. A key design choice in Bach et al. (2022) is the use of Jinja2 as a templating language, providing the flexibility crucial for crafting clear and effective prompts. Specifically, a template is a function that maps dataset examples into input-output natural language pairs, while a prompt is the combination of an input template and a target template along with a collection of specific meta-data. Instructions are specific directives within input-templates that guide the model’s behavior. However, as the original version of PromptSource did not support Persian, we modified its source code to accommodate Persian datasets. Since this system is originally integrated with Huggingface Datasets (Lhoest et al., 2021) library, we gathered datasets from various sources and consolidated into a unified public repository on HuggingFace [^3] https://huggingface.co/PNLPhub. Appendix C provides some exmaples of prompt template.
In addition to manual templating, we have decided to translate a subset of three question-answering tasks from the P3 dataset Sanh et al. (2022). This decision was made to enhance the comprehensiveness and utility of our work by providing a broader scope of data. To ensure a high-quality translation, we utilized the No Language Left Behind (NLLB) (Costa-jussà et al., 2022) machine translation model, capable of single-sentence translations between 200 languages and dialects in various scripts. We employed the largest NLLB model with 3.3B parameters to achieve the best performance. A complete list of manually templated and translated datasets is given in Figure 2.
Finally, since the datasets were sourced from multiple repositories, we applied a series of pre-processing steps such as deduplication and stripping out non-alphanumeric characters like emojis to ensure normalized text across all data.
3.2 Task Augmentation
It is widely recognized that instruction-tuned models benefit significantly from extensive and varied tasks. Given this context, we focus on developing diverse prompts, spanning from basic to proficient language levels. Furthermore, drawing from the methodologies outlined in FLAN Collection (Longpre et al., 2023), T0 (Sanh et al., 2022), and MetaICL (Min et al., 2022), we enhance task diversity by mixing and swapping different data fields within a given dataset. For instance, whereas a dataset might initially be structured to evaluate a model’s ability to answer question x given input y, we train the model to generate question x when provided with answer y. This approach effectively broadens the spectrum of prompts within a limited data pool.
3.3 Data statistics
The statistics of final dataset after applying templates across all datasets is presented in Figure 3. Table 1 also presents the total number of categorization and generation prompts for each task type.
Figure 3: Distribution of NLP tasks across the FarsInstruct dataset, highlighting the expanded data volumes post-prompt application and the number of prompts designed per task type.
Task Type Cat Gen
Question Answering 1 9
Translation 2 10
NER (Named Entity Recognition) 4 19
Multiple Choice QA 9 1
Word Sense Disambiguation 10 0
Classification 15 12
Summarization 4 15
Reading Comprehension 2 18
Query Paraphrasing 10 7
Sentiment Analysis 24 13
Textual Entailment 16 5
Table 1: List of task types, along with the number of categorization and generation prompts dedicated to each task type. The expanded version of this table can be found in the Appendix B.
3.4 Quality Control
We selectively chose publicly available Persian datasets predominantly used for single-task fine-tuning, as their extensive use ensures high quality. Furthermore, to ensure the accuracy and quality of the instructions, we conduct human evaluations through consultations with the general public and experts in the field of literature. This review process allowed us to assess the instructions from multiple perspectives and incorporate cultural and linguistic nuances, critical for ensuring the prompts’ clarity, accuracy, and relevance.
4 Methodology and Experimental Setup
To maintain our model’s robustness and generalization capabilities, we integrate the CoLA framework (Xia et al., 2024) with continual learning (Kirkpatrick et al., 2017). This section offers a thorough overview of the training procedure and evaluation setup.
4.1 Training Procedure
Given the significant computational demands of full fine-tuning, we aim to employ LoRA for the training procedure, specifically using the FarsInstruct dataset. However, as highlighted in studies by (Wang et al., 2023; Li et al., 2024), LoRA tends to underperform in multi-task training scenarios due to its limitations in capturing complex interactions between tasks, leading to suboptimal performance. To mitigate this challenge, Chain of LoRA (CoLA) (Xia et al., 2024), presents an iterative optimization framework based on the principles of the Frank-Wolfe algorithm (also known as the Conditional Gradient Method). This method involves an iterative process of fine-tuning on a single task, merging it with the base model, and reinitializing with a new LoRA module. Xia et al. (2024) shows that this process allows the model to learn higher-rank adaptations more effectively. Another persistent challenge affecting the performance of LoRA-tuned models is catastrophic forgetting. Kalajdzievski (2024) observed a strong inverse linear relationship between the fine-tuning performance and the amount of forgetting when fine-tuning LLMs with LoRA.
In this study we propose Continual-Chain of LoRA (Co-CoLA), an extension of CoLA framework which incorporates rehearsal with replay during training. More specifically, rehearsal training is an approach within the continual learning framework that involves revisiting a portion of previously learned tasks during training new tasks. Despite the limited success of continual learning frameworks, the study by Scialom et al. (2022) demonstrated that continual training of language models, such as T0 Sanh et al. (2022) with rehearsal, can effectively help them in comprehending new instruction via instruction composition, resulting in better generalization and improved performance on new tasks.
The core mathematical operation in LoRA involves updating the low-rank matrices $A$ and $B$, which are applied to modify the transformer layers of the model. The update rule can be expressed as $W^{\prime}=W+BA$ where $W$ represents the transformer layer’s original weights, and $W^{\prime}$ shows the updated weights after applying the low-rank adjustments $A$ and $B$. Essentially, Co-CoLA structures this training procedure into an iterative three phases:
Tuning: In this phase, following the standard LoRA, the base model weights remain frozen, while only the model’s LoRA parameters (represented by matrices $A$ and $B$) are fine-tuned. Additionally, a subset of previously trained data is replayed along with the new data. Formally, given the sequence $T=(T_{1},\ldots,T_{n})$ where $T_{i}$ represents the training data after applying an individual template, the training data augmented with rehearsal is defined as:
$T_{i}^{r}=T_{i}\cup\left(\sum_{j=1}^{i-1}rT_{j}\right)$ (1)
where r is the rehearsal hyper-parameter that controls the percentage of examples sampled from previous templates $T_{1},\ldots,T_{n}$.
Merging: After the tuning phase, the newly updated LoRA parameters are merged with the existing model weights. These merged weights are fixed and do not receive any gradient update in subsequent steps.
Expanding: The final phase involves preparing the model for subsequent training rounds by reinitializing the LoRA modules with new parameters ($A^{\prime}$ and $B^{\prime}$). Following Hu et al. (2021) $A^{\prime}$ adopts Gaussian initialization and $B^{\prime}$ is initialized to zero.
An illustration of this iterative three-staged approach is provided in Figure 4.
Figure 4: The Continual-Chain of LoRA Training Procedure
4.2 Evaluation Setup
Evaluation Tasks: Our model’s performance was evaluated through two categories of task types: those that were part of the training dataset ("Held in") and those introduced to the model for the first time during evaluation ("Held out"). The evaluation dataset encompasses three distinct types of tasks: Sentiment analysis and Query paraphrasing, classified as “Held in” tasks and Textual Entailment which is categorized as a “Held out” task. As illustrated in Figure 2, the evaluation includes one dataset each for sentiment analysis and paraphrase identification, alongside two datasets dedicated to entailment tasks.
Evaluation Metric: To assess the performance of our model relative to several baseline models, we utilized the ROUGE-L metric, which measures the overlap of n-grams between the generated text and reference texts. Specifically, we concentrated on the F1-scores of ROUGE-L, a metric that integrates precision and recall to provide a balanced evaluation. As demonstrated by Wang et al. (2022), the rankings produced by this metric exhibit a strong correlation with accuracy for categorization templates.
5 Results
To investigate the applicability of FarsInstruct, we choose the Ava model and instruction-tune it using the Co-CoLA framework on a diverse set of templates. Our results were compared against a series of monolingual and multilingual instruction-tuned models and to effectively assess the performance of our model we conduct both quantitative and linguistic evaluations. For a comprehensive overview of the training configuration, please refer to the Appendix A.
Task Type Model ROUGE-L
parsinlu query paraphrasing Held In Aya-13B 45.58
PersianMind-7B 17.07
Mistral-7B 6.89
Dorna-8B 3.85
Ava-8B 6.67
Ava-LoRA-8B 8.73
Co-CoLA-8B 45.86
Digikala Sentiment Analysis Held In Aya-13B 28.41
PersianMind-7B 18.19
Mistral-7B 2.46
Dorna-8B 2.42
Ava-8B 8.69
Ava-LoRA-8B 5.72
Co-CoLA-8B 40.87
FarsTail Held Out Aya-13B 37.61
PersianMind-7B 17.05
Mistral-7B 5.74
Dorna-8B 4.81
Ava-8B 12.48
Ava-LoRA-8B 9.07
Co-CoLA-8B 36.35
Parsinlu Entailment Held Out Aya-13B 42.64
PersianMind-7B 4.45
Mistral-7B 4.93
Dorna-8B 3.32
Ava-8B 15.04
Ava-LoRA-8B 7.18
Co-CoLA-8B 55.32
Table 2: ROUGE-L F1 Scores for Different Models across Tasks
5.1 Quantitative Evaluation
We evaluate our model against several existing models fine-tuned on instruction-specific data. Specifically, PersianMind (University of Tehran, 2024) is a Llama-2 7B based model, trained in 3 phases on different Persian datasets. Though their training data is unavailable, Dorna (PartAI, 2024) and Ava (Moghadam, 2024) are newly introduced models, fine-tuned on the Llama-3 8B model for Persian tasks. Aya (CohereForAI, 2024) is a 13B encoder-decoder model trained on a subset of 25 million samples from the Aya dataset and Mistral-7B (MistralAI, 2024) is a decoder-only model trained on publicly available prompted datasets
Table 2 summarizes the comparative performance of various models, including our proposed method, Co-CoLA, across several NLP Datasets: ParsiNLU Query Paraphrasing, Digikala Sentiment Analysis, FarsTail, and ParsiNLU Entailment. These models are evaluated using ROUGE-L F1 scores. As illustrated in Table 2, Co-CoLA performs comparably well to the Aya model, despite having fewer parameters and being trained on less instruction data and significantly outperforms all other models, indicating the effectiveness of Co-CoLA. The factors contributing to this performance gap are further discussed in Section 6. Moreover, the scores of Ava-LoRA, reflecting the performance of raw LoRA fine-tuning of Ava on FarsInstruct, are inferior to those achieved with Co-CoLA training, highlighting the effectiveness of our method.
Figure 5: Comparative performance of different models on Persian language tasks using the ROUGE-L metric. The bar chart depicts the superior performance of Co-CoLA across multiple tasks, particularly excelling in the ParsiNLU Entailment task.
5.2 Linguistic Evaluation
Our comprehensive linguistic evaluation aimed to further substantiate the effectiveness of Co-CoLA in handling the nuances of the Persian language, compared to the baseline model Ava. The evaluation specifically focused on analyzing the models’ capabilities in terms of coherence, relevance, and linguistic quality, which are critical for assessing the applicability of language models in real-world scenarios.
5.2.1 Evaluation Setup
The evaluation involved detailed analysis by language experts who assessed the output from both models based on predefined criteria. This approach ensures an unbiased evaluation of the models’ performance in generating contextually appropriate and linguistically accurate content.
5.2.2 Evaluation Criteria
The linguistic outputs were evaluated based on three main criteria:
Coherence: This assesses the logical flow and connectivity of the text produced by the models.
Relevance: This measures how well the model’s output adheres to the context provided in the input.
Linguistic Quality: This evaluates the grammatical accuracy, punctuation, and stylistic appropriateness of the text.
5.2.3 Evaluation Results
The evaluation results are summarized in Table 3, which provides a clear comparative analysis of the performance of the two models across all assessed criteria. The scores indicate that while Ava scored slightly higher in coherence, Co-CoLA outperformed Ava in relevance and linguistic quality, suggesting its superior ability to produce contextually accurate and linguistically refined outputs.
Criteria Co-CoLA Ava
Coherence 4.2 4.3
Relevance 3.7 3.2
Linguistic Quality 4.6 4.0
Table 3: Average Scores from Linguistic Evaluation
The higher scores of Co-CoLA in relevance and linguistic quality demonstrate its effectiveness in producing not only grammatically correct but also contextually relevant outputs, which is essential for real-world applications. These results underscore the potential of Co-CoLA in enhancing the linguistic handling of Persian language tasks, setting a benchmark for future developments in language model applications.
6 Discussion
Figure 5 provides a detailed breakdown of the overall performance reported in Table 2. Each dot in the plot represents the ROUGE-L F1 score of the given model on the selected template. As clearly illustrated, other Persian instruction-tuned models fail to achieve a high ROUGE-L F1 score. One significant factor contributing to this disparity is the low precision score. The F1 score, which combines precision and recall, serves as a comprehensive metric for evaluation. Precision measures the proportion of the longest common subsequence (LCS) in the candidate text that matches the reference text, while recall measures the proportion of the LCS in the reference text that is present in the candidate text. Although these models achieve acceptable recall scores, they fall short in precision, a critical metric for categorization templates. In contrast, Aya demonstrates proficiency in handling both generation and categorization templates within the Persian context. Compared to Aya, Co-CoLA enhances the model’s ability to manage both categorization and generation tasks effectively while being less computationally expensive.
7 Conclusion
This study introduces significant advancements with FarsInstruct and Co-CoLA, addressing critical gaps in the processing and instruction-following capabilities for Persian, a low-resource language. FarsInstruct, with its diverse tasks ranging from text summarization to named entity recognition, has proven to enhance language model performance as shown through rigorous ROUGE evaluations and human assessments. This dataset not only enriches multilingual model training but also establishes a new standard for language model instruction tuning.
Further, Co-CoLA leverages the strengths of CoLA with rehearsal training to mitigate catastrophic forgetting and improve multi-task adaptation, through its iterative optimization framework. This allows for sustained model performance over diverse tasks while optimizing computational resources. Looking ahead, the focus will be on expanding the scope of these datasets to cover more tasks and modalities, thereby driving further innovations in cross-lingual language understanding and promoting AI inclusivity.
8 Limitations
This section delineates the principal limitations of our study, which, while providing substantial contributions to Persian NLP, presents challenges that could be addressed in future developments to enhance its utility and applicability in broader linguistic contexts:
Data Diversity and Representation: Although FarsInstruct broadens the corpus of Persian language resources, it does not fully capture the rich tapestry of dialects and sociolects that characterize the Persian-speaking world. Also, the collected templates are generally biased towards short responses, which might affect the overall performance of the model.
Complexity of Instructions: The dataset prompts vary in complexity but still may not sufficiently challenge or train models to handle the types of complex instructions encountered in everyday human interactions. Real-world applications often demand a high level of interpretative depth and context awareness—qualities that current models may struggle with when trained on existing datasets. Future versions of FarsInstruct could benefit from integrating prompts that require higher-order cognitive processing, such as irony, metaphor understanding, and techniques that involve prompting the model to break down complex tasks into intermediate steps, mimicking human reasoning processes (Wei et al., 2022b).
Dependency on External Datasets: The effectiveness of the FarsInstruct dataset is contingent upon the quality and variety of the external datasets. This dependency creates vulnerability, as biases or errors in source datasets may be passed to FarsInstruct.A rigorous process for source data, coupled with efforts to develop original, high-quality training materials, could diminish reliance on external datasets and enhance the overall integrity of the dataset.
Evaluation Metrics: The metrics currently used to evaluate models trained on FarsInstruct may not fully capture the nuanced and multifaceted aspects of language comprehension and generation. Furthermore, for certain tasks such as rewriting, ROUGE-L may not serve as an adequate measure of quality.
Performance Stability: While Co-CoLA has demonstrated effectiveness in terms of computational efficiency and consistent performance across all tasks it learned, mitigating catastrophic forgetting, we observe that its overall performance is heavily dependent on the model’s performance at each tuning iteration. We leave potential solutions to this problem to future work.
Appendix
  1. A.
    Training Configuration
    All implementations were carried out using PyTorch, Transformers (Wolf et al., 2020) and Accelerate (Gugger et al., 2022) library. For efficient training, we randomly selected 25 prompt templates and applied them to their corresponding datasets. Consequently, for example, a dataset with two selected templates would be upsampled to twice its original size. We then sampled a minimum of 10,000 examples from each dataset, based on the specific template and dataset length, to create the current training data. The rehearsal hyper-parameter of Co-CoLA was set to 0.01. We used Paged-AdamW as the base optimizer and trained for a total of four epochs in each tuning phase. A linear learning rate scheduler was applied, with an initial learning rate of $6\times 10^{-5}$ and a batch size of 16. For implementing LoRA, we utilized the PEFT (Mangrulkar et al., 2022) library for convenience.
  2. B.
    Dataset Illustrations
    Dataset Categorization Generation
    DigiMag 9 1
    Digikala_sentiment_analysis 9 1
    ExaPPC 3 4
    FarsTail 8 2
    ParsABSA 5 1
    ParsiNLU_entailment 8 3
    ParsiNLU_multiple_choice 9 1
    ParsiNLU_query_paraphrasing 7 3
    ParsiNLU_reading_comprehension 1 9
    ParsiNLU_sentiment 3 7
    ParsiNLU_translation_En_FA 1 5
    ParsiNLU_translation_FA_En 1 5
    PEYMA 1 9
    Persian_NER 3 10
    Persian_news 3 3
    Persian_QA 1 9
    Pn_summary 3 8
    Snappfood_sentiment_analysis 7 4
    Syntran_FA 1 9
    Wiki_summary 1 7
    XL_WiC 10 0
    Table 4: Detailed Overview of Datasets Utilized for Categorization and Generation Tasks. Each dataset is hyperlinked to the corresponding HuggingFace repository. As shown in this table Categorization and Generation tasks are not equally distributed across all datasets. Some datasets, such as Digimag, are originally designed for categorization tasks. We have enhanced these datasets by incorporating generation prompts. Conversely, translation tasks, which are inherently generative, have been augmented with categorization prompts. This dual-purpose approach enriches the datasets, facilitating both categorization and generation tasks and providing a more versatile training and testing framework. This table provides insight into the distribution and specialization of prompts across different datasets, highlighting the balance and focus within the training and testing framework.
    Figure 6: A treemap visualization that organizes datasets by task type, post-instruction application size, and data category (training vs. testing). Each primary rectangle represents a distinct task type within the natural language processing field, encompassing areas such as Question Answering, Classification, Translation, and more. Within these primary rectangles, smaller sub-rectangles represent individual datasets. The area of each sub-rectangle is scaled to the logarithm of the size of the dataset to accommodate the broad variance in dataset sizes, ensuring a more balanced visual representation that allows for the inclusion of both large and small datasets on the same scale.
  3. C.
    Prompts
    Figure 7: A prompt example shown in Promptsource environment. PromptSource is an advanced toolkit designed for creating, sharing, and utilizing natural language prompts. Prompts function as mappings that convert examples from datasets into natural language inputs and corresponding target outputs. In PromptSource, we develop input templates, target templates, and choice templates. Inputs typically consist of questions or instructions, while the output code specifies the expected answer or result. For Categorization tasks, the choice template includes predefined options for answering questions, while Generation tasks do not require this template. In this picture, The "Metrics" box is set to measure Accuracy for Categorization tasks, and the "Prompt Language" used is Farsi (Persian). "Answer choices" are provided within the template, which comprises an instruction followed by data fields. The premise and hypothesis are selected from the "Data Schema" on the left side of the interface. The ||||||| | | symbol separates instructions from outputs, and the output employs Jinja code for conditional logic: if the label is c, it outputs (no); if the label is e, it outputs (yes); and if the label is n, it outputs (cannot determine).
    See pages 2- of figures/FarsInstruct.pdf

### v3

@@
 Abstract
-Instruction-tuned large language models, such as T0, have demonstrated remarkable capabilities in following instructions across various domains.
+Instruction-tuned large language models have demonstrated remarkable capabilities in following human instructions across various domains.
 However, their proficiency remains notably deficient in many low-resource languages.
-To address this challenge, we introduce FarsInstruct: a comprehensive instruction dataset designed to enhance the instruction-following ability of large language models specifically for the Persian language—a significant yet underrepresented language globally.
-FarsInstruct encompasses a wide range of task types and datasets, each containing a mix of straightforward to complex manual written instructions, as well as translations from Public Pool of Prompts, ensuring a rich linguistic and cultural representation.
+To address this challenge, we begin by introducing FarsInstruct: a comprehensive instruction dataset designed to enhance the instruction-following ability of large language models specifically for the Persian language—a significant yet underrepresented language globally.
+FarsInstruct encompasses a wide range of task types and datasets, each containing a mix of straightforward to complex manual written instructions, as well as translations from the Public Pool of Prompts, ensuring a rich linguistic and cultural representation.
 Furthermore, we introduce Co-CoLA, a framework designed to enhance the multi-task adaptability of LoRA-tuned models.
-Through extensive experimental analyses, our study showcases the effectiveness of FarsInstruct dataset coupled with training by Co-CoLA framework [^1] https://github.com/Hojjat-Mokhtarabadi/FarsInstruct, in improving the performance of large language models within the Persian context.
-As of the current writing, FarsInstruct comprises more than 200 templates across 21 distinct datasets, and we intend to update it consistently, thus augmenting its applicability.
+Through extensive experimental analyses, our study showcases the effectiveness of the FarsInstruct dataset coupled with training by the Co-CoLA framework, in improving the performance of large language models within the Persian context.
+As of the current writing, FarsInstruct comprises 197 templates across 21 distinct datasets, and we intend to update it consistently, thus augmenting its applicability.
+[^1] https://huggingface.co/datasets/PNLPhub/FarsInstruct
 Keywords: Instruction-tuned LLMs, Low-resource languages, Parameter efficient fine-tuning
-FarsInstruct: Empowering Large Language Models
-for Persian Instruction Understanding
-Hojjat Mokhtarabadithanks: All the authors are members of the Department of Computer Engineering.
-h.mokhtarabadi@ec.iut.ac.ir Isfahan University of Technology                        Ziba Zamanithanks: Zibazamani@eng.uk.ac.ir Shahid Bahonar University of Kerman
-Abbas Maazallahithanks: a.maazallahi@ut.ac.ir University of Tehran                        Hossein Manshaeithanks: manshaei@iut.ac.ir Isfahan University of Technology
 1 Introduction
-The modern era of artificial intelligence is marked by numerous breakthroughs, among which is the rise of large language models (LLMs).
-These models, such as PaLM (Chowdhery et al., 2022), GPT4 (OpenAI et al., 2024), and Llama2 (Touvron et al., 2023) with continuous scaling of their parameters and training data, are known to exhibit emergent properties.
-Wei et al.
-(2022a) considers an ability to be emergent if it is not present in smaller models but is present in larger models.
-This is an unpredictable phenomenon that can not be predicted simply by extrapolating the performance of smaller models.
-One such ability is instruction-following, which enables models to execute unseen natural language processing (NLP) tasks from reading instructions provided within the input text.
-Previously, the capability for instruction-following was primarily attributed to the scale of these models.
-However, recent studies have demonstrated that instruction-following does not exclusively rely on the large size of language models (Sanh et al., 2022).
-By instruction-tuning on a collection of instructional NLP tasks, smaller language models can learn to follow prompts.
-This approach has proven to be remarkably efficient, allowing these smaller models to perform competitively and, in some specific tasks, even outperform their larger counterparts (Sanh et al., 2022; Wei et al., 2021; Wang et al., 2022).
+The modern era of artificial intelligence is marked by numerous breakthroughs, among which is the rise of large language models (LLMs), such as GPT4 (openai2024gpt4), Llama3 (dubey2024llama3herdmodels) and PaLM (chowdhery2022palm).
 Instruction-tuning emerges as a vital technique in the evolution of language models, involving training a model on a wide range of tasks described through natural language instructions.
-This method diverges from traditional task-specific fine-tuning, offering a more generalized and versatile approach to model training, thus contributing significantly to the advancement of LLMs.
-Despite the steady progress of instruction-tuned language models, they still struggle to accurately grasp the subtleties of low-resourced languages due to the scarcity of prompted data and challenges inherent in translating English datasets into other languages (Naous et al., 2024; Ramesh et al., 2023; Vanmassenhove et al., 2021).
-While efforts have been made to compile extensive multilingual instruction-following datasets, gaps remain in creating diverse and complex prompts for languages like Persian.
-For example, the SuperNaturalInstructions benchmark (Wang et al., 2022), encompassing various task types across 55 languages, contains merely 2.1% of Persian content.
-Similarly, the Aya Dataset (Singh et al., 2024), a human-curated effort to enhance AI’s instruction-following abilities across 65 languages, includes 1% of Persian content.
-This underscores the disparity in the diversity and quantity of the Persian Language tasks compared to other languages.
-In this study, we propose FarsInstruct, a comprehensive prompted dataset tailored to the Persian language.
-It comprises a mixture of manually written instructions ranging from basic to proficient language levels, as well as translations from Public Pool of Prompts (P3) (Sanh et al., 2022) which is a collection of prompted English datasets.
-In particular, we created more than 200 prompt templates (roughly 10 templates for each of the 21 unique public datasets) that we selected from a variety of sources.
-These datasets collectively cover ten different task categories: Text Summarization, Textual Entailment, Text Classification, Sentiment Analysis, Word Sense Disambiguation, Query Paraphrasing, Question Answering, Reading Comprehension, Named Entity Recognition (NER), and Translation.
-Figure 1 depicts an instance of a prompt within our dataset, and a detailed overview of FarsInsturct dataset is provided in Section 3.
+This method diverges from traditional task-specific fine-tuning and adapts the model’s behavior to respond to user queries with relevant and helpful answers.
+This technique offers a more generalized and versatile approach to model training, thus contributing significantly to the advancement of LLMs.
+Despite the steady progress of instruction-tuned language models, a persistent limitation remains: their difficulty in capturing the nuanced complexities of low-resource languages.
+This critical challenge stems from the significant gap in the availability of high-quality instruction datasets tailored to these languages.
+wang2023far highlights this concern, demonstrating that datasets lacking sufficient multilingual diversity can cause models to lose previously learned multilingual capabilities, leading to performance degradation.
+Moreover, translating English-centric datasets offers only partial solutions due to several inherent limitations (naous2024having; ramesh-etal-2023-fairness; vanmassenhove-etal-2021-machine).
+While efforts have been made to compile extensive multilingual instruction datasets (wang-etal-2022-super; singh2024aya; muennighoff2022crosslingual), gaps remain in creating diverse and complex prompts for languages like Persian compared to other languages.
+In this study, we propose FarsInstruct, a comprehensive human-annotated instruction dataset created from existing Persian NLP datasets.
+It includes a mixture of manually written instructions ranging from basic to proficient language levels, alongside translations from the Public Pool of Prompts (P3) (sanh2022multitask), which is a collection of prompted English datasets.
+To ensure the diversity and representativeness of FarsInstruct, we developed 197 prompt templates derived from 21 distinct public datasets.
+Each prompt template comprises an input template and a target template, both of which function to extract relevant data fields from their respective datasets and reformat them into a unified structure designed for the instruction-tuning objective.
+For example, in the case of a Textual Entailment dataset containing the fields Premise, Hypothesis, and Label, an input template might be: ”Can the hypothesis be concluded from the premise?
+Premise: {Premise}, Hypothesis: {Hypothesis}”, while a corresponding target template could be ”The answer is: {label}”.
+The collected public datasets encompass ten different task categories: Text Summarization, Textual Entailment, Text Classification, Sentiment Analysis, Word Sense Disambiguation, Query Paraphrasing, Question Answering, Reading Comprehension, Named Entity Recognition (NER), and Translation.
+Figure 1 depicts an instance of a prompt within our dataset after applying its respective template.
+A detailed overview of the FarsInsturct dataset is provided in Section LABEL:sec:farsinstruct.
 Figure 1: An example of the prompts utilized in the training process.
 The Persian version of the prompt is employed for training purposes, while the translated English version is provided to enhance comprehension.
-The instruction component is highlighted in blue, the data field is marked in orange, and the target answer is indicated in gray.
-In Appendix C, this example is shown in PromptSource environment.
-Additionally, in order to facilitate the multi-task adaptation of our model and mitigate the problem of catastrophic forgetting, we introduce Co-CoLA, an integration of CoLA (Xia et al., 2024) with rehearsal training (Kirkpatrick et al., 2017).
+The instruction component is highlighted in black, the data fields are marked in orange, and the target answer is indicated in gray.
+In Appendix LABEL:sec:appendix_prompts, this example is shown in the PromptSource environment.
+Additionally, parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) (hu2021lora), not only face challenges in multi-task settings but are also prone to catastrophic forgetting (wang2023multilora; li2024mixlora; kalajdzievski2024scaling).
+To address these issues, we propose Co-CoLA, a novel integration of CoLA (xia2024chain) with rehearsal training (kirkpatrick2017overcoming).
 More specifically, we adopt an iterative optimization framework that merges learned low-rank matrices into the model parameters and reinitializes optimization for new LoRA modules.
-At each iteration, we involve retraining a subset of data from previously learned tasks and mixing it with the current task’s data during training.
-With this periodic revisiting of earlier tasks, the model maintains performance on both old and new tasks while preserving computational efficiency.
-Section 4 presents an in-depth explanation of Co-CoLA method.
-FarsInstruct is publicly available and open-source and we are committed to enhancing it by continually expanding our dataset with a broader range of tasks, instruction entries, and modalities.
-We hope this dataset fills the critical gap and serves as a valuable resource to the NLP community.
+At each iteration, we retrain a subset of data from previously learned tasks, mixing it with the current task’s data during training.
+This periodic revisiting of earlier tasks ensures that the model retains performance across both old and new tasks, all while preserving computational efficiency.
+Section LABEL:sec:method presents an in-depth explanation of the Co-CoLA method.
+In summary, our contributions to advancing Persian instruction understanding are threefold: (1) We present FarsInstruct, a comprehensive human-annotated instruction dataset for Persian, covering varied and representative tasks for different categories such as text summarization, named entity recognition, and translation.
+(2) We introduce Co-CoLA, a method that combines CoLA with rehearsal training to mitigate catastrophic forgetting in multi-task learning.
+(3) We release FarsInstruct as an open-source resource, with a commitment to its continued expansion to include a broader range of tasks and modalities.
 2 Related work
-Instruction-tuning: In the landscape of AI, the capabilities of LLMs have expanded far beyond mere text processing.
-These sophisticated models are now being fine-tuned in a practice known as instruction-tuning, where models are trained with specific input-output pairs drawn from a wide array of data sources.
-This technique enables a pre-trained LLM to produce tailored outputs based on given inputs, enhancing its versatility and effectiveness.
-FLAN (Wei et al., 2021) and T0 (Sanh et al., 2022) pioneered the exploration of instruction-tuned language models, each contributing significantly to the field.
-FLAN (Wei et al., 2021) adapted a 137-billion parameter pre-trained model, refining it with over 60 NLP datasets using natural language instructions.
-On the other hand, T0 (Sanh et al., 2022) applied instruction tuning to various T5 models across 2073 prompts from 177 datasets.
-SuperNaturalInstruction (Wang et al., 2022) further advanced the field by assembling a comprehensive benchmark featuring 1,616 expert-written NLP tasks, covering 76 unique task types, and extending support to multiple languages.
-xP3 (Muennighoff et al., 2022) expanded on P3’s groundwork (Sanh et al., 2022) by including content from 46 languages, adding new tasks like Translation and Program Synthesis that P3 had not tackled.
-In a similar expansive effort, Aya (Singh et al., 2024) emerged as a significant multilingual project, featuring an impressive collection of 513 million instances across 114 languages, achieved through a collaborative research effort that involved fluent speakers from around the world to compile and complete instructional content.
-Our dataset distinguishes itself from these collections in its depth and adaptability, especially with the inclusion of more challenging Persian tasks, offering a high level of detail not found in many multilingual efforts.
+Instruction-tuning: Instruction tuning refers to the process of training language models using specific input-output pairs derived from diverse data sources.
+This approach enhances the ability of a pre-trained LLM to interpret and respond to a wide range of human requests expressed in natural language.
+Instruction datasets used for this purpose are typically created in one of three ways: (1) manually created by researchers from existing NLP datasets (wang-etal-2022-super; wei2021finetuned), (2) synthesized by prompting proprietary models with a small, seed dataset (alpaca; wang2022self; honovich-etal-2023-unnatural), or (3) generated entirely from scratch, involving human-written prompt-response pairs (DatabricksBlog2023DollyV2; kopf2024openassistant).
+In this work, we adopt the first approach to develop FarsInstruct.
+Previous works such as FLAN (wei2021finetuned) and P3 (sanh2022multitask) have been instrumental in advancing instruction dataset creation.
+FLAN encompasses over 60 NLP datasets, while P3 features more than 2,000 prompts from 177 datasets, each significantly contributing to the field.
+SuperNaturalInstruction (wang-etal-2022-super) further advanced the field by assembling a comprehensive benchmark featuring 1,616 expert-written NLP tasks, covering 76 unique task types, and extending support to multiple languages.
+xP3 (muennighoff2022crosslingual) expanded on P3’s groundwork by including content from 46 languages, adding new tasks like Translation and Program Synthesis that P3 had not tackled.
+Similarly, Aya (singh2024aya) represents a major multilingual effort, featuring an extensive dataset of 513 million instances across 114 languages.
+This was achieved through a global collaboration involving fluent speakers who contributed instructional content.
+Our dataset distinguishes itself from these collections in its depth and adaptability, especially with the inclusion of more challenging tasks in Persian, offering a high level of detail not found in many multilingual efforts.
 While most such projects primarily use machine translations and cover a narrow range of tasks, our dataset presents a wide array of culturally and linguistically rich tasks.
-Parameter effecient fine-tuning: Conventional full-parameter fine-tuning becomes computationally impractical as the model size and the number of downstream tasks increase.
-To address this challenge, recent advancements in parameter-efficient fine-tuning methods suggest training only a small portion of parameters while keeping the majority of pre-trained model parameters unchanged.
-One of the most widely used paradigms in parameter-efficient fine turning is Low-Rank Adaptation (LoRA) (Hu et al., 2021).
-LoRA only modifies a small, low-rank portion of the model’s weights.
-This is achieved by adding low-rank matrices to the model’s weights during training.
-Despite the significant computational advantage of LoRA, it falls short in multi-task adaptation, and also Kalajdzievski (2024) showed that PEFT strategies, such as LoRA, are still susceptible to catastrophic forgetting.
-MultiLoRA (Wang et al., 2023) addresses the limitations of LoRA by reducing the dominance of top singular vectors, horizontally scaling LoRA modules, and altering the initialization of adaptation matrices, which leads to improved performance across multiple tasks with minimal additional parameters.
-MixLoRA (Li et al., 2024) introduces multiple LoRA-based experts within a frozen pre-trained model using a top-k routing strategy to efficiently distribute tasks, independently configure attention layer adapters, and apply auxiliary load balance loss, significantly enhancing performance while reducing GPU memory consumption and training latency.
-Additionally, CoLA(Xia et al., 2024) introduces an iterative optimization framework designed to improve the fine-tuning of LLMs by employing multiple iterations of LoRA.
+Parameter effecient fine-tuning: Conventional full-parameter fine-tuning becomes computationally impractical as model size and the number of downstream tasks increase.
+To address this challenge, recent advancements in PEFT methods advocate for training only a small subset of parameters while leaving the majority of pre-trained model parameters intact.
+One of the most widely utilized paradigms in PEFT is Low-Rank Adaptation (LoRA) (hu2021lora).
+LoRA modifies only a small, low-rank portion of the model’s weights by incorporating low-rank matrices into the model’s weights during the training process.
+Despite the significant computational advantage of LoRA, it falls short in multi-task adaptation, Additionally, kalajdzievski2024scaling demonstrated that PEFT techniques, including LoRA, remain vulnerable to catastrophic forgetting, where models lose previously acquired knowledge when fine-tuned on new tasks.
+MultiLoRA (wang2023multilora) addresses the limitations of LoRA by reducing the dominance of top singular vectors, horizontally scaling LoRA modules, and altering the initialization of adaptation matrices, which leads to improved performance across multiple tasks with minimal additional parameters.
+MixLoRA (li2024mixlora) introduces multiple LoRA-based experts within a frozen pre-trained model using a top-k routing strategy to efficiently distribute tasks, independently configure attention layer adapters, and apply auxiliary load balance loss, significantly enhancing performance while reducing GPU memory consumption and training latency.
+Further, CoLA(xia2024chain) introduces an iterative optimization framework designed to improve the fine-tuning of LLMs by employing multiple iterations of LoRA.
 In this paper, we design Co-CoLA to address the issue of catastrophic forgetting, while ensuring an effective multi-task adaption.
@@
 With about 130 million [^2] https://en.wikipedia.org/wiki/Persian_language speakers, Persian — also referred to as Farsi in Iran — is an important language in the Middle East and Central Asia.
-FarsInstruct represents a project to provide a comprehensive public prompted dataset for the Persian community.
-As of this writing, FarsInstruct has more than 200 carefully designed and created prompt templates for 21 already-published public datasets and some translations from existing prompted datasets.
-Unlike multilingual collections focusing on common tasks such as Text Summarization and Question Answering, FarsInstruct introduces more innovative and challenging tasks, including Named Entity Recognition and Word Sense Disambiguation.
+FarsInstruct represents a project to provide a comprehensive public instruction dataset for the Persian community.
+As of this writing, FarsInstruct has 197 carefully designed and created prompt templates for 21 already-published public datasets and some translations from existing prompted datasets.
+Unlike multilingual collections focusing on common tasks such as Text Summarization and Question Answering, FarsInstruct introduces more task types, including Named Entity Recognition and Word Sense Disambiguation.
 The creation procedure, statistics, task augmentation, and quality of the dataset are covered in detail in the following subsections.
-Additional illustrations and tables are provided in the Appendix BC.
+Additional illustrations and tables are provided in the Appendix LABEL:dataset_detailsLABEL:sec:appendix_dataset_assetsLABEL:sec:appendix_prompts.
 3.1 Dataset Construction
@@
 These instructions also include scenarios where the model needs to generate missing content from partial text inputs.
-Figure 2: Detailed depiction of 11 task types utilized in our dataset.
+Figure 2: The detailed depiction of 11 task types utilized in our dataset.
 Each box within the figure lists the specific datasets associated with the respective task type.
@@
 In contrast, datasets that have been translated from English to Persian are enclosed with dashed borders.
-To efficiently create a large collection of prompts, we primarily utilized PromptSource (Bach et al., 2022), an open-source tool designed for creating, sharing, and managing prompts for NLP tasks.
-A key design choice in Bach et al.
-(2022) is the use of Jinja2 as a templating language, providing the flexibility crucial for crafting clear and effective prompts.
-Specifically, a template is a function that maps dataset examples into input-output natural language pairs, while a prompt is the combination of an input template and a target template along with a collection of specific meta-data.
-Instructions are specific directives within input-templates that guide the model’s behavior.
-However, as the original version of PromptSource did not support Persian, we modified its source code to accommodate Persian datasets.
-Since this system is originally integrated with Huggingface Datasets (Lhoest et al., 2021) library, we gathered datasets from various sources and consolidated into a unified public repository on HuggingFace [^3] https://huggingface.co/PNLPhub.
-Appendix C provides some exmaples of prompt template.
-In addition to manual templating, we have decided to translate a subset of three question-answering tasks from the P3 dataset Sanh et al.
-(2022).
+To efficiently create a large collection of prompts, we primarily utilized PromptSource (bach2022promptsource), an open-source tool designed for creating, sharing, and managing prompts for NLP tasks.
+A key design choice in bach2022promptsource is the use of Jinja2333https://jinja.palletsprojects.com/en/3.1.x/ as a templating language, providing the flexibility crucial for crafting clear and effective prompts.
+Each dataset has multiple prompt templates, each of which consists of an input and a target template.
+These templates map raw data fields into natural language, structuring both the input and target sequences.
+Practically, templates allow users to mix arbitrary text with data fields.
+We refer to the text within the input template that guides the model’s behavior as ”Instruction”.
+Additionally, each prompt template documents essential metadata, including evaluation metrics and the language used.
+The PromptSource toolkit offers an interface for interactively writing prompts on datasets.
+However, the original version did not support Persian, so we modified its source code to handle Persian datasets.
+Our updated version is publicly available, providing the Persian community with a tool to simplify prompt creation  [^4] https://github.com/Hojjat-Mokhtarabadi/promptsource.
+Appendix LABEL:sec:appendix_prompts depicts an illustration of the PromptSource interface with an example of a Textual Entailment dataset.
+Moreover, since this system was originally integrated with Huggingface Datasets library (lhoest-etal-2021-datasets), we gathered datasets from various sources and consolidated them into a unified public repository on HuggingFace.
+Appendix LABEL:sec:appendix_prompts provides a sample of the crafted prompt templates for different datasets.
+In addition to manual templating, we have decided to translate a subset of three question-answering datasets from the P3 collection sanh2022multitask.
 This decision was made to enhance the comprehensiveness and utility of our work by providing a broader scope of data.
-To ensure a high-quality translation, we utilized the No Language Left Behind (NLLB) (Costa-jussà et al., 2022) machine translation model, capable of single-sentence translations between 200 languages and dialects in various scripts.
+To ensure a high-quality translation, we utilized the No Language Left Behind (NLLB) (costa2022no) machine translation model, capable of single-sentence translations between 200 languages and dialects in various scripts.
 We employed the largest NLLB model with 3.3B parameters to achieve the best performance.
 A complete list of manually templated and translated datasets is given in Figure 2.
-Finally, since the datasets were sourced from multiple repositories, we applied a series of pre-processing steps such as deduplication and stripping out non-alphanumeric characters like emojis to ensure normalized text across all data.
-3.2 Task Augmentation
-It is widely recognized that instruction-tuned models benefit significantly from extensive and varied tasks.
-Given this context, we focus on developing diverse prompts, spanning from basic to proficient language levels.
-Furthermore, drawing from the methodologies outlined in FLAN Collection (Longpre et al., 2023), T0 (Sanh et al., 2022), and MetaICL (Min et al., 2022), we enhance task diversity by mixing and swapping different data fields within a given dataset.
-For instance, whereas a dataset might initially be structured to evaluate a model’s ability to answer question x given input y, we train the model to generate question x when provided with answer y.
-This approach effectively broadens the spectrum of prompts within a limited data pool.
-3.3 Data statistics
-The statistics of final dataset after applying templates across all datasets is presented in Figure 3.
-Table 1 also presents the total number of categorization and generation prompts for each task type.
-Figure 3: Distribution of NLP tasks across the FarsInstruct dataset, highlighting the expanded data volumes post-prompt application and the number of prompts designed per task type.
+The final dataset is standardized through a series of preprocessing steps like deduplication and removing irrelevant elements (HTML tags, hyperlinks, emojis, and offensive language).
+Figure 3 shows the distribution of tasks across FarsInstruct, with Table 1 listing the total number of categorization and generation prompts for each task type.
+3.2 Task Augmentation and Quality Control
+Instruction-tuned models are widely recognized for their significant benefits from exposure to extensive and varied tasks.
+Given this context, diverse prompts encompassing a range of instructions—spanning from basic to advanced language levels—are essential.
+Building on best practices outlined in the FLAN Collection (longpre2023flan), T0 (sanh2022multitask), and MetaICL (min-etal-2022-metaicl), we enhance task diversity by mixing and swapping different data fields within a given dataset.
+For example, while a dataset may initially assess a model’s ability to answer question X based on input Y, we train the model to generate question X when provided with answer Y.
+This approach effectively broadens the range of prompts available within a limited data pool.
+Figure 3: Distribution of NLP tasks across the FarsInstruct dataset, highlighting the expanded data volumes after applying prompt templates and the number of prompts designed per task type.
+For each dataset, the final size is determined by multiplying the number of samples (N) by the number of prompt templates (M), resulting in a dataset size of N*M.
 Task Type Cat Gen
@@
 Textual Entailment 16 5
-Table 1: List of task types, along with the number of categorization and generation prompts dedicated to each task type. The expanded version of this table can be found in the Appendix B.
-3.4 Quality Control
-We selectively chose publicly available Persian datasets predominantly used for single-task fine-tuning, as their extensive use ensures high quality.
-Furthermore, to ensure the accuracy and quality of the instructions, we conduct human evaluations through consultations with the general public and experts in the field of literature.
-This review process allowed us to assess the instructions from multiple perspectives and incorporate cultural and linguistic nuances, critical for ensuring the prompts’ clarity, accuracy, and relevance.
+Table 1: List of task types, along with the number of categorization and generation prompts dedicated to each task type.
+The expanded version of this table can be found in the Appendix LABEL:sec:appendix_dataset_assets.
+To ensure the accuracy and cultural relevance of the instructions, we incorporated public input and expert evaluations.
+Feedback was gathered from 15 randomly selected individuals and three experts in Persian literature and psychology.
+Participants were asked to help craft instructions in various writing formats, including formal and informal styles, and to express the same instruction in different ways, then two psychology experts and one literature professor were consulted to refine the instructions.
+Their expertise informed revisions, ensuring that the responses were grammatically and linguistically correct and resonated with the general Persian-speaking population.
+Further, the datasets adopted in FarsInstruct are predominantly used for single-task fine-tuning, as their widespread use indicates higher quality.
 4 Methodology and Experimental Setup
-To maintain our model’s robustness and generalization capabilities, we integrate the CoLA framework (Xia et al., 2024) with continual learning (Kirkpatrick et al., 2017).
+To maintain our model’s robustness and generalization capabilities, we integrate the CoLA framework (xia2024chain) with continual learning (kirkpatrick2017overcoming).
 This section offers a thorough overview of the training procedure and evaluation setup.
@@
 Given the significant computational demands of full fine-tuning, we aim to employ LoRA for the training procedure, specifically using the FarsInstruct dataset.
-However, as highlighted in studies by (Wang et al., 2023; Li et al., 2024), LoRA tends to underperform in multi-task training scenarios due to its limitations in capturing complex interactions between tasks, leading to suboptimal performance.
-To mitigate this challenge, Chain of LoRA (CoLA) (Xia et al., 2024), presents an iterative optimization framework based on the principles of the Frank-Wolfe algorithm (also known as the Conditional Gradient Method).
+However, as highlighted in studies by (wang2023multilora; li2024mixlora), LoRA tends to underperform in multi-task training scenarios due to its limitations in capturing complex interactions between tasks, leading to suboptimal performance.
+To mitigate this challenge, Chain of LoRA (CoLA) (xia2024chain), presents an iterative optimization framework based on the principles of the Frank-Wolfe algorithm (also known as the Conditional Gradient Method).
 This method involves an iterative process of fine-tuning on a single task, merging it with the base model, and reinitializing with a new LoRA module.
-Xia et al.
-(2024) shows that this process allows the model to learn higher-rank adaptations more effectively.
+xia2024chain shows that this process allows the model to learn higher-rank adaptations more effectively.
 Another persistent challenge affecting the performance of LoRA-tuned models is catastrophic forgetting.
-Kalajdzievski (2024) observed a strong inverse linear relationship between the fine-tuning performance and the amount of forgetting when fine-tuning LLMs with LoRA.
-In this study we propose Continual-Chain of LoRA (Co-CoLA), an extension of CoLA framework which incorporates rehearsal with replay during training.
-More specifically, rehearsal training is an approach within the continual learning framework that involves revisiting a portion of previously learned tasks during training new tasks.
-Despite the limited success of continual learning frameworks, the study by Scialom et al.
-(2022) demonstrated that continual training of language models, such as T0 Sanh et al.
-(2022) with rehearsal, can effectively help them in comprehending new instruction via instruction composition, resulting in better generalization and improved performance on new tasks.
+kalajdzievski2024scaling observed a strong inverse linear relationship between the fine-tuning performance and the amount of forgetting when fine-tuning LLMs with LoRA.
+In this study, we propose Continual-Chain of LoRA (Co-CoLA), an extension of the CoLA framework that incorporates rehearsal with replay during training.
+More specifically, rehearsal training is an approach within the continual learning framework that involves revisiting a portion of previously learned tasks while training new tasks.
 The core mathematical operation in LoRA involves updating the low-rank matrices $A$ and $B$, which are applied to modify the transformer layers of the model.
 The update rule can be expressed as $W^{\prime}=W+BA$ where $W$ represents the transformer layer’s original weights, and $W^{\prime}$ shows the updated weights after applying the low-rank adjustments $A$ and $B$.
-Essentially, Co-CoLA structures this training procedure into an iterative three phases:
-Tuning: In this phase, following the standard LoRA, the base model weights remain frozen, while only the model’s LoRA parameters (represented by matrices $A$ and $B$) are fine-tuned.
-Additionally, a subset of previously trained data is replayed along with the new data.
-Formally, given the sequence $T=(T_{1},\ldots,T_{n})$ where $T_{i}$ represents the training data after applying an individual template, the training data augmented with rehearsal is defined as:
-$T_{i}^{r}=T_{i}\cup\left(\sum_{j=1}^{i-1}rT_{j}\right)$ (1)
-where r is the rehearsal hyper-parameter that controls the percentage of examples sampled from previous templates $T_{1},\ldots,T_{n}$.
-Merging: After the tuning phase, the newly updated LoRA parameters are merged with the existing model weights.
+Essentially, Co-CoLA structures the training procedure by iterating over the following three phases:
+Tuning: Following the standard LoRA approach, the weights of the base model remain frozen, while only the model’s LoRA parameters, represented by matrices $A$ and $B$ are fine-tuned.
+During this phase, a subset of previously trained data is replayed along with the new data.
+Formally, let $T=(T_{1},\ldots,T_{n})$ denote the sequence where each $T_{i}$ represents the training data obtained by applying the prompt template $i$ to its corresponding dataset.
+the training data augmented with rehearsal is defined as:
+$T_{i}^{r}=T_{i}\cup\left(\sum_{j=1}^{i-1}rT_{j}\right)$
+where r is the rehearsal hyperparameter that controls the percentage of examples sampled from previous tasks $T_{1},\ldots,T_{i-1}$.
+Merging: After the tuning phase, the newly updated LoRA parameters are merged with the existing model weights based on the standard method in hu2021lora.
 These merged weights are fixed and do not receive any gradient update in subsequent steps.
 Expanding: The final phase involves preparing the model for subsequent training rounds by reinitializing the LoRA modules with new parameters ($A^{\prime}$ and $B^{\prime}$).
-Following Hu et al.
-(2021) $A^{\prime}$ adopts Gaussian initialization and $B^{\prime}$ is initialized to zero.
+Following hu2021lora $A^{\prime}$ adopts Gaussian initialization and $B^{\prime}$ is initialized to zero.
 An illustration of this iterative three-staged approach is provided in Figure 4.
@@
 4.2 Evaluation Setup
-Evaluation Tasks: Our model’s performance was evaluated through two categories of task types: those that were part of the training dataset ("Held in") and those introduced to the model for the first time during evaluation ("Held out").
-The evaluation dataset encompasses three distinct types of tasks: Sentiment analysis and Query paraphrasing, classified as “Held in” tasks and Textual Entailment which is categorized as a “Held out” task.
-As illustrated in Figure 2, the evaluation includes one dataset each for sentiment analysis and paraphrase identification, alongside two datasets dedicated to entailment tasks.
-Evaluation Metric: To assess the performance of our model relative to several baseline models, we utilized the ROUGE-L metric, which measures the overlap of n-grams between the generated text and reference texts.
-Specifically, we concentrated on the F1-scores of ROUGE-L, a metric that integrates precision and recall to provide a balanced evaluation.
-As demonstrated by Wang et al.
-(2022), the rankings produced by this metric exhibit a strong correlation with accuracy for categorization templates.
+The performance of our model is assessed across two categories of task types: those included in the training dataset (”Held in”) and those introduced for the first time during evaluation (”Held out”).
+This choice allows for a more comprehensive evaluation of the model’s generalization abilities.
+The evaluation dataset comprises three distinct task types: Sentiment Analysis and Query Paraphrasing, classified as “Held in” tasks, and Textual Entailment, categorized as a “Held out” task.
+As shown in Figure 2, the evaluation includes one dataset each for sentiment analysis and paraphrase identification, as well as two datasets specifically for entailment tasks.
+We employ the ROUGE-L metric to evaluate the overlap of n-grams between the generated text and reference texts.
+Our focus was on the F1-scores of ROUGE-L, which combines precision and recall for a comprehensive assessment.
+As shown by wang-etal-2022-super, the rankings generated by this metric correlate strongly with accuracy for categorization templates.
 5 Results
-To investigate the applicability of FarsInstruct, we choose the Ava model and instruction-tune it using the Co-CoLA framework on a diverse set of templates.
-Our results were compared against a series of monolingual and multilingual instruction-tuned models and to effectively assess the performance of our model we conduct both quantitative and linguistic evaluations.
-For a comprehensive overview of the training configuration, please refer to the Appendix A.
-Task Type Model ROUGE-L
-parsinlu query paraphrasing Held In Aya-13B 45.58
-PersianMind-7B 17.07
-Mistral-7B 6.89
-Dorna-8B 3.85
-Ava-8B 6.67
-Ava-LoRA-8B 8.73
-Co-CoLA-8B 45.86
-Digikala Sentiment Analysis Held In Aya-13B 28.41
-PersianMind-7B 18.19
-Mistral-7B 2.46
-Dorna-8B 2.42
-Ava-8B 8.69
-Ava-LoRA-8B 5.72
-Co-CoLA-8B 40.87
-FarsTail Held Out Aya-13B 37.61
-PersianMind-7B 17.05
-Mistral-7B 5.74
-Dorna-8B 4.81
-Ava-8B 12.48
-Ava-LoRA-8B 9.07
-Co-CoLA-8B 36.35
-Parsinlu Entailment Held Out Aya-13B 42.64
-PersianMind-7B 4.45
-Mistral-7B 4.93
-Dorna-8B 3.32
-Ava-8B 15.04
-Ava-LoRA-8B 7.18
-Co-CoLA-8B 55.32
-Table 2: ROUGE-L F1 Scores for Different Models across Tasks
-5.1 Quantitative Evaluation
-We evaluate our model against several existing models fine-tuned on instruction-specific data.
-Specifically, PersianMind (University of Tehran, 2024) is a Llama-2 7B based model, trained in 3 phases on different Persian datasets.
-Though their training data is unavailable, Dorna (PartAI, 2024) and Ava (Moghadam, 2024) are newly introduced models, fine-tuned on the Llama-3 8B model for Persian tasks.
-Aya (CohereForAI, 2024) is a 13B encoder-decoder model trained on a subset of 25 million samples from the Aya dataset and Mistral-7B (MistralAI, 2024) is a decoder-only model trained on publicly available prompted datasets
-Table 2 summarizes the comparative performance of various models, including our proposed method, Co-CoLA, across several NLP Datasets: ParsiNLU Query Paraphrasing, Digikala Sentiment Analysis, FarsTail, and ParsiNLU Entailment.
-These models are evaluated using ROUGE-L F1 scores.
-As illustrated in Table 2, Co-CoLA performs comparably well to the Aya model, despite having fewer parameters and being trained on less instruction data and significantly outperforms all other models, indicating the effectiveness of Co-CoLA.
-The factors contributing to this performance gap are further discussed in Section 6.
-Moreover, the scores of Ava-LoRA, reflecting the performance of raw LoRA fine-tuning of Ava on FarsInstruct, are inferior to those achieved with Co-CoLA training, highlighting the effectiveness of our method.
-Figure 5: Comparative performance of different models on Persian language tasks using the ROUGE-L metric.
-The bar chart depicts the superior performance of Co-CoLA across multiple tasks, particularly excelling in the ParsiNLU Entailment task.
-5.2 Linguistic Evaluation
-Our comprehensive linguistic evaluation aimed to further substantiate the effectiveness of Co-CoLA in handling the nuances of the Persian language, compared to the baseline model Ava.
-The evaluation specifically focused on analyzing the models’ capabilities in terms of coherence, relevance, and linguistic quality, which are critical for assessing the applicability of language models in real-world scenarios.
-5.2.1 Evaluation Setup
-The evaluation involved detailed analysis by language experts who assessed the output from both models based on predefined criteria.
-This approach ensures an unbiased evaluation of the models’ performance in generating contextually appropriate and linguistically accurate content.
-5.2.2 Evaluation Criteria
-The linguistic outputs were evaluated based on three main criteria:
-Coherence: This assesses the logical flow and connectivity of the text produced by the models.
-Relevance: This measures how well the model’s output adheres to the context provided in the input.
-Linguistic Quality: This evaluates the grammatical accuracy, punctuation, and stylistic appropriateness of the text.
-5.2.3 Evaluation Results
-The evaluation results are summarized in Table 3, which provides a clear comparative analysis of the performance of the two models across all assessed criteria.
-The scores indicate that while Ava scored slightly higher in coherence, Co-CoLA outperformed Ava in relevance and linguistic quality, suggesting its superior ability to produce contextually accurate and linguistically refined outputs.
-Criteria Co-CoLA Ava
-Coherence 4.2 4.3
-Relevance 3.7 3.2
-Linguistic Quality 4.6 4.0
-Table 3: Average Scores from Linguistic Evaluation
-The higher scores of Co-CoLA in relevance and linguistic quality demonstrate its effectiveness in producing not only grammatically correct but also contextually relevant outputs, which is essential for real-world applications.
-These results underscore the potential of Co-CoLA in enhancing the linguistic handling of Persian language tasks, setting a benchmark for future developments in language model applications.
-6 Discussion
-Figure 5 provides a detailed breakdown of the overall performance reported in Table 2.
-Each dot in the plot represents the ROUGE-L F1 score of the given model on the selected template.
-As clearly illustrated, other Persian instruction-tuned models fail to achieve a high ROUGE-L F1 score.
-One significant factor contributing to this disparity is the low precision score.
-The F1 score, which combines precision and recall, serves as a comprehensive metric for evaluation.
-Precision measures the proportion of the longest common subsequence (LCS) in the candidate text that matches the reference text, while recall measures the proportion of the LCS in the reference text that is present in the candidate text.
-Although these models achieve acceptable recall scores, they fall short in precision, a critical metric for categorization templates.
-In contrast, Aya demonstrates proficiency in handling both generation and categorization templates within the Persian context.
-Compared to Aya, Co-CoLA enhances the model’s ability to manage both categorization and generation tasks effectively while being less computationally expensive.
+To investigate whether FarsInstruct can enhance LLM’s performance on Persian tasks, we perform instruction tuning on an open-source Persian-English LLM as the foundation model.
+We additionally employ mistral, aya and llama as baselines.
+In our experiments, we discovered that the FarsInstruct dataset significantly boosts the instruction-following capabilities of language models tailored for the Persian language.
+Initial results indicate that models trained with FarsInstruct not only improved in traditional NLP tasks like text classification and summarization but also showed enhanced performance in complex tasks such as entailment and sentiment analysis.
+Specifically, the use of FarsInstruct resulted in an average improvement of 15% in accuracy scores across various models when compared to those trained with existing datasets.
+This enhancement underscores the efficacy of FarsInstruct in addressing the gaps in data diversity and complexity for low-resource languages.
+Moreover, the models exhibited increased robustness in handling nuanced linguistic features of Persian, a critical aspect often overlooked in standard datasets.
+During the evaluation phase, models demonstrated an adeptness in interpreting intricate instructions and generating responses that align more closely with the expectations set by native Persian speakers.
+The success of FarsInstruct in elevating the performance of Persian language models offers promising avenues for future research in multilingual NLP, particularly for languages that are underrepresented in the digital space.
 6 Conclusion
-This study introduces significant advancements with FarsInstruct and Co-CoLA, addressing critical gaps in the processing and instruction-following capabilities for Persian, a low-resource language.
-FarsInstruct, with its diverse tasks ranging from text summarization to named entity recognition, has proven to enhance language model performance as shown through rigorous ROUGE evaluations and human assessments.
-This dataset not only enriches multilingual model training but also establishes a new standard for language model instruction tuning.
-Further, Co-CoLA leverages the strengths of CoLA with rehearsal training to mitigate catastrophic forgetting and improve multi-task adaptation, through its iterative optimization framework.
-This allows for sustained model performance over diverse tasks while optimizing computational resources.
-Looking ahead, the focus will be on expanding the scope of these datasets to cover more tasks and modalities, thereby driving further innovations in cross-lingual language understanding and promoting AI inclusivity.
-8 Limitations
-This section delineates the principal limitations of our study, which, while providing substantial contributions to Persian NLP, presents challenges that could be addressed in future developments to enhance its utility and applicability in broader linguistic contexts:
-Data Diversity and Representation: Although FarsInstruct broadens the corpus of Persian language resources, it does not fully capture the rich tapestry of dialects and sociolects that characterize the Persian-speaking world.
-Also, the collected templates are generally biased towards short responses, which might affect the overall performance of the model.
-Complexity of Instructions: The dataset prompts vary in complexity but still may not sufficiently challenge or train models to handle the types of complex instructions encountered in everyday human interactions.
-Real-world applications often demand a high level of interpretative depth and context awareness—qualities that current models may struggle with when trained on existing datasets.
-Future versions of FarsInstruct could benefit from integrating prompts that require higher-order cognitive processing, such as irony, metaphor understanding, and techniques that involve prompting the model to break down complex tasks into intermediate steps, mimicking human reasoning processes (Wei et al., 2022b).
-Dependency on External Datasets: The effectiveness of the FarsInstruct dataset is contingent upon the quality and variety of the external datasets.
-This dependency creates vulnerability, as biases or errors in source datasets may be passed to FarsInstruct.A rigorous process for source data, coupled with efforts to develop original, high-quality training materials, could diminish reliance on external datasets and enhance the overall integrity of the dataset.
-Evaluation Metrics: The metrics currently used to evaluate models trained on FarsInstruct may not fully capture the nuanced and multifaceted aspects of language comprehension and generation.
-Furthermore, for certain tasks such as rewriting, ROUGE-L may not serve as an adequate measure of quality.
-Performance Stability: While Co-CoLA has demonstrated effectiveness in terms of computational efficiency and consistent performance across all tasks it learned, mitigating catastrophic forgetting, we observe that its overall performance is heavily dependent on the model’s performance at each tuning iteration.
-We leave potential solutions to this problem to future work.
-Appendix
-  1. A.
-    Training Configuration
-    All implementations were carried out using PyTorch, Transformers (Wolf et al., 2020) and Accelerate (Gugger et al., 2022) library.
-For efficient training, we randomly selected 25 prompt templates and applied them to their corresponding datasets.
-Consequently, for example, a dataset with two selected templates would be upsampled to twice its original size.
-We then sampled a minimum of 10,000 examples from each dataset, based on the specific template and dataset length, to create the current training data.
-The rehearsal hyper-parameter of Co-CoLA was set to 0.01.
-We used Paged-AdamW as the base optimizer and trained for a total of four epochs in each tuning phase.
-A linear learning rate scheduler was applied, with an initial learning rate of $6\times 10^{-5}$ and a batch size of 16.
-For implementing LoRA, we utilized the PEFT (Mangrulkar et al., 2022) library for convenience.
-  2. B.
-    Dataset Illustrations
+In this work we introduce farsinstruct, a comprehensive instruction dataset specifically curated for Persian language,[bridging the critical gap in current language resources for underrepresented language and advancing LLMs training in cross lingual domain] to bridge the critical gap [sth] [what gap?].
+Farsinstruct is comprised of 210 templates across 11 distinct datasets.
+Also to ensure the richeness of data we translated 3 popular question answering datasets.
+Looking ahead, we are committed to the ongoing enrichment and refinement of Farsinstruct.
+We will incorporate a wider array of task types, instruction entries, and modalities, in line with the latest advances in cross lingual language understanding.
+Moreover given the limitations imposed by Lora training strategy and computational resources current language models have yet to improve on Persian context.
+V2
+In concluding the study on ”FarsInstruct: Empowering Large Language Models for Persian Instruction Understanding,” it becomes evident that this groundbreaking dataset marks a pivotal advancement in the field of NLP for low-resource languages, particularly Persian.
+By meticulously curating over 200 templates from a variety of sources and enriching the linguistic landscape with complex Persian tasks, FarsInstruct not only bridges a critical gap in multilingual model training but also sets a new precedent for future research in language model instruction tuning.
+The dataset’s comprehensive coverage, reflecting a rich array of task types and linguistic nuances, promises significant improvements in model performance and instruction following capabilities.
+Looking ahead, the commitment to continual updates and expansion of FarsInstruct underscores the ongoing quest for achieving greater linguistic diversity and model adaptability.
+This endeavor not only enriches the Persian NLP toolkit but also inspires further exploration and innovation in the domain of instruction-tuned large language models, potentially catalyzing advancements in cross-lingual understanding and global inclusivity in AI technologies.
+Acknowledgments
+Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
+Appendix A Appendix
+A.1
 Dataset Categorization Generation
@@
 The area of each sub-rectangle is scaled to the logarithm of the size of the dataset to accommodate the broad variance in dataset sizes, ensuring a more balanced visual representation that allows for the inclusion of both large and small datasets on the same scale.
-  3. C.
-    Prompts
-    Figure 7: A prompt example shown in Promptsource environment.
+A.2 Prompts
+Figure 6: An example shown in the PromptSource environment.
 PromptSource is an advanced toolkit designed for creating, sharing, and utilizing natural language prompts.
-Prompts function as mappings that convert examples from datasets into natural language inputs and corresponding target outputs.
+Prompt templates function as mappings that convert examples from datasets into natural language inputs and corresponding target outputs.
 In PromptSource, we develop input templates, target templates, and choice templates.
@@
 For categorization tasks, the choice template includes predefined options for answering questions, while generation tasks do not require this template.
-In this picture, The "Metrics" box is set to measure Accuracy for Categorization tasks, and the "Prompt Language" used is Farsi (Persian).
-"Answer choices" are provided within the template, which comprises an instruction followed by data fields.
-The premise and hypothesis are selected from the "Data Schema" on the left side of the interface.
+In this picture, The ”Metrics” box is set to measure Accuracy for categorization tasks, and the ”Prompt Language” used is Farsi (Persian).
+”Answer choices” are provided within the template, which comprises an instruction followed by data fields.
+The premise and hypothesis are selected from the ”Data Schema” on the left side of the interface.
 The ||||||| | | symbol separates instructions from outputs, and the output employs Jinja code for conditional logic: if the label is c, it outputs (no); if the label is e, it outputs (yes); and if the label is n, it outputs (cannot determine).

### v4

@@
-FarsInstruct: Empowering Large Language Models for Persian Instruction Understanding
+Empowering Persian LLMs for Instruction Following:
 Abstract
@@
 As of the current writing, FarsInstruct comprises 197 templates across 21 distinct datasets, and we intend to update it consistently, thus augmenting its applicability.
-[^1] https://huggingface.co/datasets/PNLPhub/FarsInstruct
 Keywords: Instruction-tuned LLMs, Low-resource languages, Parameter efficient fine-tuning
+Empowering Persian LLMs for Instruction Following:
+A Novel Dataset and Training Approachthanks: Accepted at LoResLM @ COLING 2025
+Hojjat Mokhtarabadi, Ziba Zamani, Abbas Maazallahi, Mohammad Hossein Manshaei§ Department of Electrical and Computer Engineering, Isfahan University of Technology, Iran Department of Computer Engineering, Shahid Bahonar University of Kerman, Iran Department of Electrical and Computer Engineering, University of Tehran, Tehran, Iran §Department of Computer Science, Hunter College, City University of New York, NY, USA h.mokhtarabadi@ec.iut.ac.ir, zibazamani@eng.uk.ac.ir, a.maazallahi@ut.ac.ir, mohammad.manshaei@hunter.cuny.edu
 1 Introduction
-The modern era of artificial intelligence is marked by numerous breakthroughs, among which is the rise of large language models (LLMs), such as GPT4 (openai2024gpt4), Llama3 (dubey2024llama3herdmodels) and PaLM (chowdhery2022palm).
+The modern era of artificial intelligence is marked by numerous breakthroughs, among which is the rise of large language models (LLMs), such as GPT4 (OpenAI et al., 2024), Llama3 (Dubey et al., 2024) and PaLM (Chowdhery et al., 2022).
 Instruction-tuning emerges as a vital technique in the evolution of language models, involving training a model on a wide range of tasks described through natural language instructions.
@@
 This critical challenge stems from the significant gap in the availability of high-quality instruction datasets tailored to these languages.
-wang2023far highlights this concern, demonstrating that datasets lacking sufficient multilingual diversity can cause models to lose previously learned multilingual capabilities, leading to performance degradation.
-Moreover, translating English-centric datasets offers only partial solutions due to several inherent limitations (naous2024having; ramesh-etal-2023-fairness; vanmassenhove-etal-2021-machine).
-While efforts have been made to compile extensive multilingual instruction datasets (wang-etal-2022-super; singh2024aya; muennighoff2022crosslingual), gaps remain in creating diverse and complex prompts for languages like Persian compared to other languages.
+Wang et al.
+(2023b) highlights this concern, demonstrating that datasets lacking sufficient multilingual diversity can cause models to lose previously learned multilingual capabilities, leading to performance degradation.
+Moreover, translating English-centric datasets offers only partial solutions due to several inherent limitations (Naous et al., 2024; Ramesh et al., 2023; Vanmassenhove et al., 2021).
+While efforts have been made to compile extensive multilingual instruction datasets (Wang et al., 2022b; Singh et al., 2024; Muennighoff et al., 2022), gaps remain in creating diverse and complex prompts for languages like Persian compared to other languages.
 In this study, we propose FarsInstruct, a comprehensive human-annotated instruction dataset created from existing Persian NLP datasets.
-It includes a mixture of manually written instructions ranging from basic to proficient language levels, alongside translations from the Public Pool of Prompts (P3) (sanh2022multitask), which is a collection of prompted English datasets.
+It includes a mixture of manually written instructions ranging from basic to proficient language levels, alongside translations from the Public Pool of Prompts (P3) (Sanh et al., 2022), which is a collection of prompted English datasets.
 To ensure the diversity and representativeness of FarsInstruct, we developed 197 prompt templates derived from 21 distinct public datasets.
 Each prompt template comprises an input template and a target template, both of which function to extract relevant data fields from their respective datasets and reformat them into a unified structure designed for the instruction-tuning objective.
-For example, in the case of a Textual Entailment dataset containing the fields Premise, Hypothesis, and Label, an input template might be: ”Can the hypothesis be concluded from the premise?
-Premise: {Premise}, Hypothesis: {Hypothesis}”, while a corresponding target template could be ”The answer is: {label}”.
+For example, in the case of a Textual Entailment dataset containing the fields Premise, Hypothesis, and Label, an input template might be: "Can the hypothesis be concluded from the premise?
+Premise: {Premise}, Hypothesis: {Hypothesis}", while a corresponding target template could be "The answer is: {label}".
 The collected public datasets encompass ten different task categories: Text Summarization, Textual Entailment, Text Classification, Sentiment Analysis, Word Sense Disambiguation, Query Paraphrasing, Question Answering, Reading Comprehension, Named Entity Recognition (NER), and Translation.
 Figure 1 depicts an instance of a prompt within our dataset after applying its respective template.
-A detailed overview of the FarsInsturct dataset is provided in Section LABEL:sec:farsinstruct.
+A detailed overview of the FarsInsturct dataset is provided in Section 3.
 Figure 1: An example of the prompts utilized in the training process.
@@
 The instruction component is highlighted in black, the data fields are marked in orange, and the target answer is indicated in gray.
-In Appendix LABEL:sec:appendix_prompts, this example is shown in the PromptSource environment.
-Additionally, parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) (hu2021lora), not only face challenges in multi-task settings but are also prone to catastrophic forgetting (wang2023multilora; li2024mixlora; kalajdzievski2024scaling).
-To address these issues, we propose Co-CoLA, a novel integration of CoLA (xia2024chain) with rehearsal training (kirkpatrick2017overcoming).
+In Appendix D, this example is shown in the PromptSource environment.
+Additionally, parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) (Hu et al., 2021), not only face challenges in multi-task settings but are also prone to catastrophic forgetting (Wang et al., 2023a; Li et al., 2024; Kalajdzievski, 2024).
+To address these issues, we propose Co-CoLA, a novel integration of CoLA (Xia et al., 2024) with rehearsal training (Kirkpatrick et al., 2017).
 More specifically, we adopt an iterative optimization framework that merges learned low-rank matrices into the model parameters and reinitializes optimization for new LoRA modules.
@@
 This periodic revisiting of earlier tasks ensures that the model retains performance across both old and new tasks, all while preserving computational efficiency.
-Section LABEL:sec:method presents an in-depth explanation of the Co-CoLA method.
+Section 4 presents an in-depth explanation of the Co-CoLA method.
 In summary, our contributions to advancing Persian instruction understanding are threefold: (1) We present FarsInstruct, a comprehensive human-annotated instruction dataset for Persian, covering varied and representative tasks for different categories such as text summarization, named entity recognition, and translation.
 (2) We introduce Co-CoLA, a method that combines CoLA with rehearsal training to mitigate catastrophic forgetting in multi-task learning.
-(3) We release FarsInstruct as an open-source resource, with a commitment to its continued expansion to include a broader range of tasks and modalities.
+(3) We release FarsInstruct as an open-source resource, with a commitment to its continued expansion to include a broader range of tasks and modalities [^1] https://huggingface.co/datasets/PNLPhub/FarsInstruct, [^2] https://github.com/Hojjat-Mokhtarabadi/FarsInstruct.
 2 Related work
@@
 This approach enhances the ability of a pre-trained LLM to interpret and respond to a wide range of human requests expressed in natural language.
-Instruction datasets used for this purpose are typically created in one of three ways: (1) manually created by researchers from existing NLP datasets (wang-etal-2022-super; wei2021finetuned), (2) synthesized by prompting proprietary models with a small, seed dataset (alpaca; wang2022self; honovich-etal-2023-unnatural), or (3) generated entirely from scratch, involving human-written prompt-response pairs (DatabricksBlog2023DollyV2; kopf2024openassistant).
+Instruction datasets used for this purpose are typically created in one of three ways: (1) manually created by researchers from existing NLP datasets (Wang et al., 2022b; Wei et al., 2021), (2) synthesized by prompting proprietary models with a small, seed dataset (Taori et al., 2023; Wang et al., 2022a; Honovich et al., 2023), or (3) generated entirely from scratch, involving human-written prompt-response pairs (Conover et al., 2023; Köpf et al., 2024).
 In this work, we adopt the first approach to develop FarsInstruct.
-Previous works such as FLAN (wei2021finetuned) and P3 (sanh2022multitask) have been instrumental in advancing instruction dataset creation.
+Previous works such as FLAN (Wei et al., 2021) and P3 (Sanh et al., 2022) have been instrumental in advancing instruction dataset creation.
 FLAN encompasses over 60 NLP datasets, while P3 features more than 2,000 prompts from 177 datasets, each significantly contributing to the field.
-SuperNaturalInstruction (wang-etal-2022-super) further advanced the field by assembling a comprehensive benchmark featuring 1,616 expert-written NLP tasks, covering 76 unique task types, and extending support to multiple languages.
-xP3 (muennighoff2022crosslingual) expanded on P3’s groundwork by including content from 46 languages, adding new tasks like Translation and Program Synthesis that P3 had not tackled.
-Similarly, Aya (singh2024aya) represents a major multilingual effort, featuring an extensive dataset of 513 million instances across 114 languages.
+SuperNaturalInstruction (Wang et al., 2022b) further advanced the field by assembling a comprehensive benchmark featuring 1,616 expert-written NLP tasks, covering 76 unique task types, and extending support to multiple languages.
+xP3 (Muennighoff et al., 2022) expanded on P3’s groundwork by including content from 46 languages, adding new tasks like Translation and Program Synthesis that P3 had not tackled.
+Similarly, Aya (Singh et al., 2024) represents a major multilingual effort, featuring an extensive dataset of 513 million instances across 114 languages.
 This was achieved through a global collaboration involving fluent speakers who contributed instructional content.
@@
 To address this challenge, recent advancements in PEFT methods advocate for training only a small subset of parameters while leaving the majority of pre-trained model parameters intact.
-One of the most widely utilized paradigms in PEFT is Low-Rank Adaptation (LoRA) (hu2021lora).
+One of the most widely utilized paradigms in PEFT is Low-Rank Adaptation (LoRA) (Hu et al., 2021).
 LoRA modifies only a small, low-rank portion of the model’s weights by incorporating low-rank matrices into the model’s weights during the training process.
-Despite the significant computational advantage of LoRA, it falls short in multi-task adaptation, Additionally, kalajdzievski2024scaling demonstrated that PEFT techniques, including LoRA, remain vulnerable to catastrophic forgetting, where models lose previously acquired knowledge when fine-tuned on new tasks.
-MultiLoRA (wang2023multilora) addresses the limitations of LoRA by reducing the dominance of top singular vectors, horizontally scaling LoRA modules, and altering the initialization of adaptation matrices, which leads to improved performance across multiple tasks with minimal additional parameters.
-MixLoRA (li2024mixlora) introduces multiple LoRA-based experts within a frozen pre-trained model using a top-k routing strategy to efficiently distribute tasks, independently configure attention layer adapters, and apply auxiliary load balance loss, significantly enhancing performance while reducing GPU memory consumption and training latency.
-Further, CoLA(xia2024chain) introduces an iterative optimization framework designed to improve the fine-tuning of LLMs by employing multiple iterations of LoRA.
+Despite the significant computational advantage of LoRA, it falls short in multi-task adaptation, Additionally, Kalajdzievski (2024) demonstrated that PEFT techniques, including LoRA, remain vulnerable to catastrophic forgetting, where models lose previously acquired knowledge when fine-tuned on new tasks.
+MultiLoRA (Wang et al., 2023a) addresses the limitations of LoRA by reducing the dominance of top singular vectors, horizontally scaling LoRA modules, and altering the initialization of adaptation matrices, which leads to improved performance across multiple tasks with minimal additional parameters.
+MixLoRA (Li et al., 2024) introduces multiple LoRA-based experts within a frozen pre-trained model using a top-k routing strategy to efficiently distribute tasks, independently configure attention layer adapters, and apply auxiliary load balance loss, significantly enhancing performance while reducing GPU memory consumption and training latency.
+Further, CoLA (Xia et al., 2024) introduces an iterative optimization framework designed to improve the fine-tuning of LLMs by employing multiple iterations of LoRA.
 In this paper, we design Co-CoLA to address the issue of catastrophic forgetting, while ensuring an effective multi-task adaption.
@@
 The creation procedure, statistics, task augmentation, and quality of the dataset are covered in detail in the following subsections.
-Additional illustrations and tables are provided in the Appendix LABEL:dataset_detailsLABEL:sec:appendix_dataset_assetsLABEL:sec:appendix_prompts.
+Additional illustrations and tables are provided in the Appendix BCD.
 3.1 Dataset Construction
+Figure 2: The detailed depiction of 11 task types utilized in our dataset.
+Each box within the figure lists the specific datasets associated with the respective task type.
+Datasets designated for training are highlighted in blue, and those reserved for testing are marked in orange.
+Additionally, manual datasets, which have been specifically curated and prompted by our team, are enclosed with solid borders.
+In contrast, datasets that have been translated from English to Persian are enclosed with dashed borders.
 The development of FarsInstruct entailed transforming Persian NLP datasets into their prompted format, described in plain language.
@@
 These instructions also include scenarios where the model needs to generate missing content from partial text inputs.
-Figure 2: The detailed depiction of 11 task types utilized in our dataset.
-Each box within the figure lists the specific datasets associated with the respective task type.
-Datasets designated for training are highlighted in blue, and those reserved for testing are marked in orange.
-Additionally, manual datasets, which have been specifically curated and prompted by our team, are enclosed with solid borders.
-In contrast, datasets that have been translated from English to Persian are enclosed with dashed borders.
-To efficiently create a large collection of prompts, we primarily utilized PromptSource (bach2022promptsource), an open-source tool designed for creating, sharing, and managing prompts for NLP tasks.
-A key design choice in bach2022promptsource is the use of Jinja2333https://jinja.palletsprojects.com/en/3.1.x/ as a templating language, providing the flexibility crucial for crafting clear and effective prompts.
+To efficiently create a large collection of prompts, we primarily utilized PromptSource (Bach et al., 2022), an open-source tool designed for creating, sharing, and managing prompts for NLP tasks.
+A key design choice in Bach et al.
+(2022) is the use of Jinja2444https://jinja.palletsprojects.com/en/3.1.x/ as a templating language, providing the flexibility crucial for crafting clear and effective prompts.
 Each dataset has multiple prompt templates, each of which consists of an input and a target template.
@@
 Practically, templates allow users to mix arbitrary text with data fields.
-We refer to the text within the input template that guides the model’s behavior as ”Instruction”.
+We refer to the text within the input template that guides the model’s behavior as "Instruction".
 Additionally, each prompt template documents essential metadata, including evaluation metrics and the language used.
@@
 However, the original version did not support Persian, so we modified its source code to handle Persian datasets.
-Our updated version is publicly available, providing the Persian community with a tool to simplify prompt creation  [^4] https://github.com/Hojjat-Mokhtarabadi/promptsource.
-Appendix LABEL:sec:appendix_prompts depicts an illustration of the PromptSource interface with an example of a Textual Entailment dataset.
-Moreover, since this system was originally integrated with Huggingface Datasets library (lhoest-etal-2021-datasets), we gathered datasets from various sources and consolidated them into a unified public repository on HuggingFace.
-Appendix LABEL:sec:appendix_prompts provides a sample of the crafted prompt templates for different datasets.
-In addition to manual templating, we have decided to translate a subset of three question-answering datasets from the P3 collection sanh2022multitask.
+Our updated version is publicly available, providing the Persian community with a tool to simply create and develop prompts [^5] https://github.com/Hojjat-Mokhtarabadi/promptsource.
+Appendix D depicts an illustration of the PromptSource interface with an example of a Textual Entailment dataset.
+Moreover, since this system was originally integrated with Huggingface Datasets library (Lhoest et al., 2021), we gathered datasets from various sources and consolidated them into a unified public repository on HuggingFace.
+Appendix D provides a sample of the crafted prompt templates for different datasets.
+In addition to manual templating, we have decided to translate a subset of three question-answering datasets from the P3 collection Sanh et al.
+(2022).
 This decision was made to enhance the comprehensiveness and utility of our work by providing a broader scope of data.
-To ensure a high-quality translation, we utilized the No Language Left Behind (NLLB) (costa2022no) machine translation model, capable of single-sentence translations between 200 languages and dialects in various scripts.
+To ensure a high-quality translation, we utilized the No Language Left Behind (NLLB) (Costa-jussà et al., 2022) machine translation model, capable of single-sentence translations between 200 languages and dialects in various scripts.
 We employed the largest NLLB model with 3.3B parameters to achieve the best performance.
@@
 3.2 Task Augmentation and Quality Control
-Instruction-tuned models are widely recognized for their significant benefits from exposure to extensive and varied tasks.
-Given this context, diverse prompts encompassing a range of instructions—spanning from basic to advanced language levels—are essential.
-Building on best practices outlined in the FLAN Collection (longpre2023flan), T0 (sanh2022multitask), and MetaICL (min-etal-2022-metaicl), we enhance task diversity by mixing and swapping different data fields within a given dataset.
-For example, while a dataset may initially assess a model’s ability to answer question X based on input Y, we train the model to generate question X when provided with answer Y.
-This approach effectively broadens the range of prompts available within a limited data pool.
+Instruction-tuned language models are known for their significant benefits from exposure to a broad array of tasks.
+In this regard, we aimed to diversify the tasks through two approaches.
+First, we phrased the instructions at varying language levels, ranging from basic to advanced.
+Second, building on best practices outlined in the FLAN Collection (Longpre et al., 2023), T0 (Sanh et al., 2022), and MetaICL (Min et al., 2022), we enhanced task diversity by mixing and swapping different data fields within a given dataset.
+For instance, while a dataset may initially assess a model’s ability to answer question X based on input Y, we train the model to generate question X when provided with answer Y, thereby effectively broadening the range of prompts available within a limited data pool.
 Figure 3: Distribution of NLP tasks across the FarsInstruct dataset, highlighting the expanded data volumes after applying prompt templates and the number of prompts designed per task type.
@@
 Textual Entailment 16 5
-Table 1: List of task types, along with the number of categorization and generation prompts dedicated to each task type.
-The expanded version of this table can be found in the Appendix LABEL:sec:appendix_dataset_assets.
+Table 1: List of task types, along with the number of categorization and generation prompts dedicated to each task type. The expanded version of this table can be found in the Appendix C.
 To ensure the accuracy and cultural relevance of the instructions, we incorporated public input and expert evaluations.
@@
 4 Methodology and Experimental Setup
-To maintain our model’s robustness and generalization capabilities, we integrate the CoLA framework (xia2024chain) with continual learning (kirkpatrick2017overcoming).
+To maintain our model’s robustness and generalization capabilities, we integrate the CoLA framework (Xia et al., 2024) with continual learning (Kirkpatrick et al., 2017).
 This section offers a thorough overview of the training procedure and evaluation setup.
 4.1 Training Procedure
+Figure 4: The Continual-Chain of LoRA training procedure, containing Tuning, Merging, and Expanding.
+In Step 1, the pretrained language model is LoRA-tuned on dataset_1, with the replay memory initialized as empty and merged.
+In Step 2, the model is expanded with a new LoRA module and further tuned on a subset of dataset_1, determined by the rehearsal hyperparameter, alongside dataset_2, preparing it for Step 3.
+This process is iteratively repeated in subsequent steps.
 Given the significant computational demands of full fine-tuning, we aim to employ LoRA for the training procedure, specifically using the FarsInstruct dataset.
-However, as highlighted in studies by (wang2023multilora; li2024mixlora), LoRA tends to underperform in multi-task training scenarios due to its limitations in capturing complex interactions between tasks, leading to suboptimal performance.
-To mitigate this challenge, Chain of LoRA (CoLA) (xia2024chain), presents an iterative optimization framework based on the principles of the Frank-Wolfe algorithm (also known as the Conditional Gradient Method).
-This method involves an iterative process of fine-tuning on a single task, merging it with the base model, and reinitializing with a new LoRA module.
-xia2024chain shows that this process allows the model to learn higher-rank adaptations more effectively.
+However, as noted in the studies by (Wang et al., 2023a; Li et al., 2024), LoRA tends to underperform in multi-task training scenarios due to its limitations in capturing complex interactions between tasks, leading to suboptimal performance.
+To mitigate this challenge, Chain of LoRA (CoLA) (Xia et al., 2024), presents an iterative optimization framework based on the principles of the Frank-Wolfe algorithm (Frank et al., 1956).
+This method involves an iterative process of LoRA fine-tuning on a single task, merging the learned parameters with the base model, and reinitializing with a new LoRA module.
+Xia et al.
+(2024) shows this process allows the model to learn higher-rank adaptations more effectively.
 Another persistent challenge affecting the performance of LoRA-tuned models is catastrophic forgetting.
-kalajdzievski2024scaling observed a strong inverse linear relationship between the fine-tuning performance and the amount of forgetting when fine-tuning LLMs with LoRA.
+Kalajdzievski (2024) analyzed this phenomenon and revealed that forgetting significantly undermines both model safety and performance on reasoning benchmarks.
 In this study, we propose Continual-Chain of LoRA (Co-CoLA), an extension of the CoLA framework that incorporates rehearsal with replay during training.
@@
 The training data augmented with rehearsal is defined as:
-$T_{i}^{r}=T_{i}\cup\left(\sum_{j=1}^{i-1}rT_{j}\right)$
+$T_{i}^{r}=T_{i}\cup\left(\sum_{j=1}^{i-1}rT_{j}\right)$ (1)
 where r is the rehearsal hyperparameter that controls the percentage of examples sampled from previous tasks $T_{1},\ldots,T_{i-1}$.
-Merging: After the tuning phase, the newly updated LoRA parameters are merged with the existing model weights based on the standard method in hu2021lora.
+Merging: After the tuning phase, the newly updated LoRA parameters are merged with the existing model weights based on the standard method in Hu et al.
+(2021).
 These merged weights are fixed and do not receive any gradient update in subsequent steps.
 Expanding: The final phase involves preparing the model for subsequent training rounds by reinitializing the LoRA modules with new parameters ($A^{\prime}$ and $B^{\prime}$).
-Following hu2021lora $A^{\prime}$ adopts Gaussian initialization and $B^{\prime}$ is initialized to zero.
+Following Hu et al.
+(2021) $A^{\prime}$ adopts Gaussian initialization and $B^{\prime}$ is initialized to zero.
 An illustration of this iterative three-staged approach is provided in Figure 4.
-Figure 4: The Continual-Chain of LoRA Training Procedure
 4.2 Evaluation Setup
-The performance of our model is assessed across two categories of task types: those included in the training dataset (”Held in”) and those introduced for the first time during evaluation (”Held out”).
+The performance of our model is assessed across two categories of task types: those included in the training dataset ("Held in") and those introduced for the first time during evaluation ("Held out").
 This choice allows for a more comprehensive evaluation of the model’s generalization abilities.
@@
 Our focus was on the F1-scores of ROUGE-L, which combines precision and recall for a comprehensive assessment.
-As shown by wang-etal-2022-super, the rankings generated by this metric correlate strongly with accuracy for categorization templates.
+As shown by Wang et al.
+(2022b), the rankings generated by this metric correlate strongly with accuracy for categorization templates.
 5 Results
-To investigate whether FarsInstruct can enhance LLM’s performance on Persian tasks, we perform instruction tuning on an open-source Persian-English LLM as the foundation model.
-We additionally employ mistral, aya and llama as baselines.
-In our experiments, we discovered that the FarsInstruct dataset significantly boosts the instruction-following capabilities of language models tailored for the Persian language.
-Initial results indicate that models trained with FarsInstruct not only improved in traditional NLP tasks like text classification and summarization but also showed enhanced performance in complex tasks such as entailment and sentiment analysis.
-Specifically, the use of FarsInstruct resulted in an average improvement of 15% in accuracy scores across various models when compared to those trained with existing datasets.
-This enhancement underscores the efficacy of FarsInstruct in addressing the gaps in data diversity and complexity for low-resource languages.
-Moreover, the models exhibited increased robustness in handling nuanced linguistic features of Persian, a critical aspect often overlooked in standard datasets.
-During the evaluation phase, models demonstrated an adeptness in interpreting intricate instructions and generating responses that align more closely with the expectations set by native Persian speakers.
-The success of FarsInstruct in elevating the performance of Persian language models offers promising avenues for future research in multilingual NLP, particularly for languages that are underrepresented in the digital space.
+To investigate the applicability of FarsInstruct, we instruction-tuned Ava—a Llama-3-based Persian LLM—using the Co-CoLA framework across a suit of tasks.
+The results were compared with monolingual and multilingual instruction-tuned models, using quantitative evaluations.
+For a comprehensive overview of the training configuration, please refer to the Appendix A.
+Task Type Model ROUGE-L
+parsinlu query paraphrasing Held In Aya-13B 45.58
+PersianMind-7B 17.07
+Mistral-7B 6.89
+Ava-8B 6.67
+Ava-LoRA-8B 8.73
+CoLA-8B 20.88
+Co-CoLA-8B 45.86
+Digikala Sentiment Analysis Held In Aya-13B 28.41
+PersianMind-7B 18.19
+Mistral-7B 2.46
+Ava-8B 8.69
+Ava-LoRA-8B 5.72
+Ava-LoRA-8B 5.72
+CoLA-8B 25.62
+Co-CoLA-8B 40.87
+FarsTail Held Out Aya-13B 37.61
+PersianMind-7B 17.05
+Mistral-7B 5.74
+Ava-8B 12.48
+Ava-LoRA-8B 9.07
+CoLA-8B 15.64
+Co-CoLA-8B 36.35
+Parsinlu Entailment Held Out Aya-13B 42.64
+PersianMind-7B 4.45
+Mistral-7B 4.93
+Ava-8B 15.04
+Ava-LoRA-8B 7.18
+CoLA-8B 22.55
+Co-CoLA-8B 55.32
+Table 2: ROUGE-L F1 Scores for Different Models across Tasks
+5.1 Quantitative Evaluation
+Figure 5: Comparative performance of different models on Persian language tasks using the ROUGE-L metric.
+The bar chart depicts the superior performance of Co-CoLA across multiple tasks, particularly excelling in the ParsiNLU Entailment task.
+We evaluate our model against several existing models fine-tuned on instruction data.
+Specifically, PersianMind (University of Tehran, 2024) is a Llama-2 7B-based model, trained in 3 phases on different Persian datasets.
+Though its training data is unavailable, Ava (Moghadam, 2024) is a newly introduced model, fine-tuned on the Llama-3 8B model for Persian tasks.
+Aya (CohereForAI, 2024) is a 13B encoder-decoder model trained on a subset of 25 million samples from the Aya dataset and Mistral-7B (MistralAI, 2024) is a decoder-only model trained on publicly available prompted datasets.
+Table 2 summarizes the comparative performance of various models, including our proposed method, Co-CoLA, across several NLP Datasets: ParsiNLU Query Paraphrasing, Digikala Sentiment Analysis, FarsTail, and ParsiNLU Entailment.
+These models are evaluated using ROUGE-L F1 scores.
+As illustrated in Table 2, Co-CoLA performs comparably well to the Aya model, despite having fewer parameters and being trained on less instruction data and significantly outperforms all other models, indicating the effectiveness of Co-CoLA.
+The factors contributing to this performance gap are further discussed in Section 6.
+Moreover, the scores of Ava-LoRA, reflecting the performance of raw LoRA fine-tuning of Ava on FarsInstruct and naive CoLA are inferior to those achieved with Co-CoLA training, highlighting the effectiveness of our method.
+6 Discussion
+Figure 5 provides a detailed breakdown of the overall performance reported in Table 2.
+Each dot in the plot represents the ROUGE-L F1 score of the given model on the selected template.
+As clearly illustrated, other Persian instruction-tuned models fail to achieve a high ROUGE-L F1 score.
+One significant factor contributing to this disparity is the low precision score.
+The F1 score combines precision and recall and serves as a comprehensive metric for evaluation.
+Precision measures the proportion of the longest common subsequence (LCS) in the candidate text that matches the reference text, while recall measures the proportion of the LCS in the reference text that is present in the candidate text.
+Although these models achieve acceptable recall scores, they fall short in precision, a critical metric for categorization templates.
+In contrast, Aya demonstrates proficiency in handling both generation and categorization templates within the Persian context.
+Compared to Aya, Co-CoLA enhances the model’s ability to manage both categorization and generation tasks effectively while being less computationally expensive.
+Despite the limited success of continual learning frameworks, the study by Scialom et al.
+(2022) demonstrated that continual training of language models, such as T0 (Sanh et al., 2022) with rehearsal, can effectively help them in comprehending new instructions via instruction composition.
+Our results confirm this finding within the Chain-of-LoRA framework, resulting in better generalization and improved performance on new tasks.
 7 Conclusion
-In this work we introduce farsinstruct, a comprehensive instruction dataset specifically curated for Persian language,[bridging the critical gap in current language resources for underrepresented language and advancing LLMs training in cross lingual domain] to bridge the critical gap [sth] [what gap?].
-Farsinstruct is comprised of 210 templates across 11 distinct datasets.
-Also to ensure the richeness of data we translated 3 popular question answering datasets.
-Looking ahead, we are committed to the ongoing enrichment and refinement of Farsinstruct.
-We will incorporate a wider array of task types, instruction entries, and modalities, in line with the latest advances in cross lingual language understanding.
-Moreover given the limitations imposed by Lora training strategy and computational resources current language models have yet to improve on Persian context.
-V2
-In concluding the study on ”FarsInstruct: Empowering Large Language Models for Persian Instruction Understanding,” it becomes evident that this groundbreaking dataset marks a pivotal advancement in the field of NLP for low-resource languages, particularly Persian.
-By meticulously curating over 200 templates from a variety of sources and enriching the linguistic landscape with complex Persian tasks, FarsInstruct not only bridges a critical gap in multilingual model training but also sets a new precedent for future research in language model instruction tuning.
-The dataset’s comprehensive coverage, reflecting a rich array of task types and linguistic nuances, promises significant improvements in model performance and instruction following capabilities.
-Looking ahead, the commitment to continual updates and expansion of FarsInstruct underscores the ongoing quest for achieving greater linguistic diversity and model adaptability.
-This endeavor not only enriches the Persian NLP toolkit but also inspires further exploration and innovation in the domain of instruction-tuned large language models, potentially catalyzing advancements in cross-lingual understanding and global inclusivity in AI technologies.
-Acknowledgments
-Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
-Appendix A Appendix
-A.1
+This study aims to address the limitations in instruction-following capabilities of language models for Persian, an important but underrepresented language, by introducing a novel instruction dataset and a training approach specifically designed to enhance the instruction comprehension of large language models.
+FarsInstruct presents a carefully curated dataset that combines human-annotated instruction data with translations from English-centric instruction datasets, featuring tasks in different forms and from varying levels of difficulty.
+Further, Co-CoLA leverages the strengths of CoLA with rehearsal training to mitigate catastrophic forgetting and improve multi-task adaptation, through its iterative optimization framework.
+Our results demonstrate that this allows for sustained model performance over diverse tasks while optimizing computational resources.
+We hope our dataset fills the critical gap and serves as a valuable resource to the multilingual NLP community.
+8 Limitations
+This section delineates the principal limitations of our study, which, while providing substantial contributions to Persian NLP, presents certain challenges.
+Addressing these challenges in future developments could enhance its utility and applicability in broader linguistic contexts.
+Data Diversity and Representation: Although FarsInstruct broadens the corpus of Persian language resources, it does not fully capture the rich tapestry of dialects and sociolects that characterize the Persian-speaking world.
+Also, the collected templates are generally biased towards short responses, which might affect the overall performance of the model.
+Complexity of Instructions: The dataset prompts vary in complexity but still may not sufficiently challenge or train models to handle the types of complex instructions encountered in everyday human interactions.
+Real-world applications often demand a high level of interpretative depth and context awareness—qualities that current models may struggle with when trained on existing datasets.
+Future versions of FarsInstruct could benefit from integrating prompts that require higher-order cognitive processing, such as irony, metaphor understanding, and techniques that involve prompting the model to break down complex tasks into intermediate steps, mimicking human reasoning processes (Wei et al., 2022).
+Dependency on External Datasets: The effectiveness of the FarsInstruct dataset is contingent upon the quality and variety of the external datasets.
+This dependency creates vulnerability, as biases or errors in source datasets may be passed to FarsInstruct.
+A rigorous process for source data, coupled with efforts to develop original, high-quality training materials, could diminish reliance on external datasets and enhance the overall integrity of the dataset.
+Evaluation Metrics: The metrics currently used to evaluate models trained on FarsInstruct may not fully capture the nuanced and multifaceted aspects of language comprehension and generation.
+Furthermore, for certain tasks such as rewriting, ROUGE-L may not serve as an adequate measure of quality.
+Performance Stability: While Co-CoLA has demonstrated effectiveness in terms of computational efficiency and consistent performance across all tasks it learned, mitigating catastrophic forgetting, we observe that its overall performance is heavily dependent on the model’s performance at each tuning iteration.
+We leave potential solutions to this problem to future work.
+Appendix
+Appendix A Training Configuration
+All implementations were carried out using PyTorch Paszke et al.
+(2019), Transformers (Wolf et al., 2020) and Accelerate (Gugger et al., 2022) libraries.
+For efficient training, we randomly selected 25 prompt templates and applied them to their corresponding datasets.
+Consequently, for instance, a dataset with two selected templates would be upsampled to twice its original size.
+To generate the training data for each iteration, we sampled up to 10,000 instances from the dataset based on the selected template, with the rehearsal hyperparameter of Co-CoLA set to 0.01.
+Following the established practices, we used Paged-AdamW as the base optimizer and trained for a total of four epochs in each tuning phase.
+A linear learning rate scheduler was applied, with an initial learning rate of $6\times 10^{-5}$ and a batch size of 16.
+For implementing LoRA, we utilized the PEFT (Mangrulkar et al., 2022) library for convenience.
+Appendix B Datasets Details
+  • Digikala Sentiment Analysis (Tehranipour, 2019): A collection of Digikala product reviews labeled by customer star ratings. It categorizes sentiment into five labels (e.g., buy, not buy, neutral).
+  • Snappfood Sentiment Analysis (Tehranipour, 2022): A dataset of 70,000 user reviews from Snappfood, an online food delivery service.
+It contains equal numbers of positive and negative reviews (35,000 each), supporting effective sentiment analysis.
+  • ParsiNLU (Khashabi et al., 2021): A comprehensive suite for Persian NLP tasks, covering reading comprehension, multiple-choice question-answering, sentiment analysis, textual entailment, question-answering, and machine translation.
+These datasets are collected in a multitude of ways, often involving manual annotations by native speakers.
+This results in over 14.5k new instances across 6 distinct NLU tasks, serving as a key Persian NLP benchmark.
+  • ExaPPC (Sadeghi et al., 2022): A paraphrase corpus with 2.3 million Persian sentence pairs labeled as paraphrase or non-paraphrase.
+It includes both formal and colloquial sentences, making it ideal for models like BERT.
+  • FarsTail (Amirkhani et al., 2023): A Persian textual entailment dataset with 10,367 samples, classifying premise-hypothesis pairs into entailment, contradiction, or neutral, essential for natural language inference in Persian.
+  • Pars-ABSA (Ataei et al., 2019): A dataset for aspect-based sentiment analysis in Persian, with 5,114 positive, 3,061 negative, and 1,827 neutral data points.
+It is useful for studying fine-grained sentiment in reviews.
+  • WikiSummary Farahani (2020): A summarization dataset with 45,654 entries derived from Persian Wikipedia articles, paired with highlights designed for summarization tasks with reduced article lengths.
+  • Pn-Summary (Farahani et al., 2021): The Pn-Summary dataset contains 93,207 news articles from six news agencies, each paired with a human-generated summary.
+The dataset was curated from an initial pool of 200,000 articles, covering various categories.
+  • PersianQA (Ayoubi, 2021): PersianQA is a reading comprehension dataset with over 9,000 entries sourced from Persian Wikipedia, including both answerable and unanswerable questions.
+It supports the development of models that can recognize unanswerable queries, similar to SQuAD 2.0.
+  • PersianNews (Mehrdad Farahani, 2020): This dataset consists of 16,438 news articles from online Persian news agencies, categorized into eight classes such as Economic, International, Political, Science & Technology, and Sport.
+  • DigiMag (Mehrdad Farahani, 2020): DigiMag contains 8,515 articles from the Digikala Online Magazine, divided into seven categories including Video Games, Shopping Guide, Health & Beauty, and Art & Cinema.
+  • PEYMA (Shahshahani et al., 2018): The PEYMA dataset features 7,145 sentences with 302,530 tokens, 41,148 of which are annotated with seven entity classes, including Organization, Money, Location, Date, and Person.
+  • Persian NER (Poostchi et al., 2016): This is a manually-annotated named entity recognition dataset with 250,015 tokens and 7,682 sentences.
+The dataset includes six named entity classes like Person, Organization, Location, and Event, in IOB format.
+  • Syntran-fa (Farsi et al., 2024): A Farsi question-answering dataset with nearly 50,000 question-answer pairs.
+It extends short answers into fluent, complete responses using syntactic rules and parsing methods like stanza.
+  • XL-WiC (Raganato et al., 2020): XL-WiC is a multilingual dataset for word sense disambiguation, involving binary classification of word meaning across 12 languages, including Farsi.
+It evaluates models on cross-lingual semantic contextualization.
+  • SciQ (Lu et al., 2022): A multimodal dataset with 21,208 science questions from elementary and high school curricula.
+It covers various sciences, with questions annotated with images, lectures, and explanations, making it a rich resource for science QA.
+  • TriviaQA (Joshi et al., 2017): A large QA dataset with 950,000 question-answer pairs from Wikipedia and web documents.
+It is more challenging than datasets like SQuAD due to longer contexts and non-direct text spans, including human-verified and machine-generated pairs.
+  • AdversarialQA (Bartolo et al., 2020): A dataset designed to test the robustness of QA models against adversarially crafted questions.
+It includes adversarially modified questions from SQuAD, TriviaQA, and NewsQA to challenge model reasoning and generalization.
+Appendix C Datasets Illustrations
 Dataset Categorization Generation
@@
 Table 3: Detailed Overview of Datasets Utilized for Categorization and Generation Tasks.
-Each dataset is hyperlinked to the corresponding HuggingFace repository.
 As shown in this table Categorization and Generation tasks are not equally distributed across all datasets.
@@
 The area of each sub-rectangle is scaled to the logarithm of the size of the dataset to accommodate the broad variance in dataset sizes, ensuring a more balanced visual representation that allows for the inclusion of both large and small datasets on the same scale.
-A.2 Prompts
+Appendix D Prompts
 Figure 7: An example shown in the PromptSource environment.
@@
 For categorization tasks, the choice template includes predefined options for answering questions, while generation tasks do not require this template.
-In this picture, The ”Metrics” box is set to measure Accuracy for categorization tasks, and the ”Prompt Language” used is Farsi (Persian).
-”Answer choices” are provided within the template, which comprises an instruction followed by data fields.
-The premise and hypothesis are selected from the ”Data Schema” on the left side of the interface.
+In this picture, The "Metrics" box is set to measure Accuracy for categorization tasks, and the "Prompt Language" used is Farsi (Persian).
+"Answer choices" are provided within the template, which comprises an instruction followed by data fields.
+The premise and hypothesis are selected from the "Data Schema" on the left side of the interface.
 The ||||||| | | symbol separates instructions from outputs, and the output employs Jinja code for conditional logic: if the label is c, it outputs (no); if the label is e, it outputs (yes); and if the label is n, it outputs (cannot determine).